### v2

# THE RESTRICTED PICARD FUNCTOR

Abstract
We prove in significant generality the (almost-)representability of the Picard functor when restricted to smooth test schemes. The novelty lies in the fact that we prove such (almost-)representability beyond the proper setting.
1 Introduction
The Picard functor of a variety plays an important role in number theory and algebraic geometry. Some fundamental examples are the dual variety of an abelian variety and the Jacobian of a projective curve, both of which are defined in terms of Picard schemes. Picard schemes also arise in the construction of the Albanese variety of a proper variety, as well as in many other contexts.
Representability results for Picard varieties are therefore of fundamental importance. A wonderful theorem, proved by Murre [Mur] and Oort [Oor] building on work of Grothendieck [FGA], states that the Picard functor $\operatorname{Pic}_{X/K}$ of a proper scheme over a field $K$ is represented by a scheme [Mur, Th. 2]. Beyond the proper case one cannot expect such representability, as one already sees in the case of the affine line. Indeed, the Picard group of the affine line over a field is trivial, so were $\operatorname{Pic}_{\mathbf{A}^{1}/K}$ represented by a scheme (necessarily locally of finite type), it would have to be infinitesimal. But for a reduced scheme $X$ that is not seminormal, one does not generally have $\operatorname{Pic}(\mathbf{A}^{1}_{X})=\operatorname{Pic}(X)$, so the desired representability cannot hold. A better, more conceptual obstacle to representability is that a simple cohomological calculation using the ring of dual numbers as a test scheme shows that the tangent space to $\operatorname{Pic}_{X/K}$ at the identity is given by ${\rm{H}}^{1}(X,\mathcal{O}_{X})$, which will not be finite in general beyond the proper setting.
One observes that the problem in the above examples arises from test schemes that are not “nice.” Thus Achet, following a suggestion of Raynaud, was led to introduce the restricted Picard functor – which is the Picard functor but with one’s test schemes restricted to be smooth over $K$ – and proved its representability for forms of affine space admitting a regular compactification [Ach, Th. 1.1].
Definition 1.1.
Let $K$ be a field and $X$ a $K$-scheme. The naive restricted Picard functor
$(\operatorname{Pic}^{+}_{X/K})^{\rm{naive}}\colon\{\mbox{smooth }K-\mbox{schemes}\}\rightarrow\{\mbox{abelian groups}\}$
is the functor defined by the formula $T\mapsto\operatorname{Pic}(X\times T)$. The restricted Picard functor $\operatorname{Pic}^{+}_{X/K}$ is the étale sheafification of $(\operatorname{Pic}^{+}_{X/K})^{\rm{naive}}$.
Remark 1.2.
The étale sheafification in the definition of the restricted Picard functor is essential if one wishes to obtain representability results, as schemes are always sheaves for the étale (and even the fpqc) site, by the general theory of descent. The same technical issue arises in the definition of the usual Picard functor; see [Kle, §9.2]. For an example of a $K$-scheme $X$ such that the map $\operatorname{Pic}(X)\rightarrow\operatorname{Pic}^{+}_{X/K}(K)$ is not surjective, see Example 6.10.
Unfortunately, even the restricted Picard functor cannot be represented by a scheme in general. To illustrate the problem, let $\overline{C}/K$ be a smooth projective geometrically connected curve of genus $g>0$, let $x,y\in\overline{C}(K)$ be distinct rational points, and let $C:=\overline{C}\backslash\{x,y\}$. Then $\operatorname{Pic}_{\overline{C}/K}$ is an extension of $\mathbf{Z}$ by an abelian variety, the Jacobian $J$ of $\overline{C}$. Because $\operatorname{Pic}(\overline{C}\times T)\twoheadrightarrow\operatorname{Pic}(C\times T)$ for every smooth (or even regular) $K$-scheme $T$, one can conclude that $\operatorname{Pic}^{+}_{C/K}$ is the quotient of $J$ by the point $j$ represented by the divisor $[x]-[y]$. If $j$ is not a torsion point, then this quotient is not a scheme. Thus we see that an inevitable obstacle to representability comes from quotients of group schemes by subgroups generated by points of infinite order. We may now state the main result of the present paper. For a commutative, locally finite type $K$-group scheme $G$, we denote by $G^{t}$ the preimage in $G$ of the torsion subgroup of the étale component group $G/G^{0}$.
Theorem 1.3.
If $K$ is a field and $X$ is a regular $K$-scheme of finite type, then there is an isomorphism $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$, where $G$ is a smooth commutative $K$-group scheme with finitely generated component group and $E\subset G^{t}$ is an étale $K$-group scheme such $E(K_{s})$ is a free abelian group of finite rank. Furthermore, $G^{0}$ is unique up to unique isomorphism.
For results on the structure of the group $G^{0}$ appearing in Theorem 1.3, see §6. As we shall see, the most technically demanding case of Theorem 1.3 is when $K$ is imperfect. Indeed, for perfect $K$, one may entirely skip §§34.
Theorem 1.3 is nice, but what we would really like are results asserting representability by a scheme. Using the above theorem, however, it will not be hard to deduce the following.
Theorem 1.4.
Let $K$ be a field and $X$ a regular $K$-scheme of finite type with dense smooth locus. Assume that any map from $X_{K_{s}}$ into an abelian variety $A/K_{s}$ factors through a finite subscheme of $A$. Then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative $K$-group scheme with finitely generated component group. Furthermore, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is wound unipotent.
As examples of the above, let us mention two particular results.
Theorem 1.5.
Let $K$ be a field and let $X$ be a regular $K$-scheme of finite type with dense smooth locus. If $X_{\overline{K}}$ admits a dense open subscheme that is rationally connected, then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative $K$-group scheme with finitely generated component group. Furthermore, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is wound unipotent.
Proof.
Because abelian varieties do not admit nonconstant maps from rationally connected varieties, $X$ admits no nonconstant map to an abelian variety over $\overline{K}$, hence also not over $K_{s}$. Now apply Theorem 1.4. ∎
As an easy consequence of the above, we will prove the following theorem, which generalizes Achet’s result to arbitrary forms of affine space, with no assumption on the existence of a regular compactification. Following Achet, we call a smooth connected unipotent $K$-group $U$ strongly wound when it admits no nontrivial unirational $K$-subgroup. Equivalently, there is no nonconstant $K$-morphism $X\rightarrow U$ with $X$ an open subscheme of $\mathbf{P}^{1}_{K}$.
Theorem 1.6.
If $K$ is a field and $X$ is a $K$-form of $\mathbf{A}^{n}_{K}$, then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative unipotent $K$-group scheme with strongly wound identity component. If $X(K)\neq\emptyset$, then, for every smooth $K$-scheme $S$, $\operatorname{Pic}^{+}_{X/K}(S)=\operatorname{Pic}(X\times S)/\operatorname{Pic}(S)$.
Among the examples falling under the rubric of Theorem 1.6 are smooth connected unipotent groups. These may indeed have positive-dimensional restricted Picard functors when $K$ is imperfect. For examples, see for instance [Ros2, Props. 5.2, 5.4, 5.5, Ex. 5.8].
The proof of Theorem 1.3 proceeds roughly as follows. If $X$ admits a regular compactification $\overline{X}$ whose complementary divisors are all geometrically irreducible, then for any smooth connected $K$-scheme $T$, one has an exact sequence
$\oplus_{D}\mathbf{Z}\longrightarrow\operatorname{Pic}(\overline{X}\times T)\longrightarrow\operatorname{Pic}(X\times T)\longrightarrow 0,$
where the sum is over the divisors on $\overline{X}$ disjoint from $X$. We know that the proper $\overline{X}$ has representable Picard scheme, and we can use this and the above sequence to construct $\operatorname{Pic}^{+}_{X/K}$ as a quotient of $\operatorname{Pic}_{\overline{X}/K}$ by a finitely generated subgroup. (This is a slight simplification: Because $\operatorname{Pic}_{\overline{X}/K}$ may be nonsmooth, one actually forms the quotient of a maximal smooth subgroup scheme.) This is roughly speaking the strategy used in [Ach] to deal with forms of affine space having regular compactification.
The problem, of course, is that we do not know whether $X$ admits a regular compactification. Even when $K$ is algebraically closed, in positive characteristic the existence of a regular compactification remains a formidable open problem. Fortunately, however, there is a weaker form of resolution which is adequate for many purposes: de Jong’s alterations theorem [deJ1, Th. 4.1]. (Actually, we will require a somewhat more precise result from [deJ2].) The difference between an alteration and a resolution is that we relax the requirement that the resolution be a birational map and only require that it be generically finite. We may replace $X$ by a dense open subscheme and thereby assume that the alteration $X^{\prime}\rightarrow X$ is finite flat. We then use faithfully flat descent to describe a line bundle on $X$ in terms of descent datum for a line bundle on $X^{\prime}$. This is helpful because $X^{\prime}$ admits a regular compactification and therefore has “almost-representable” restricted Picard functor in the sense of Theorem 1.3.
The key difficulty in showing that the functor whose points consist of a line bundle plus descent datum on $X^{\prime}$ is (almost) representable is to prove that the descent data which arise in this manner are “uniformly bounded” in a suitable sense. To illustrate the difficulty, consider the situation in which $X=\mathbf{A}^{1}_{K}$ and $X^{\prime}=\mathbf{A}^{1}_{K^{\prime}}$ for some finite extension field $K^{\prime}/K$. (Of course, this is in a sense a silly example because $\operatorname{Pic}^{+}_{X/K}=0$ in this case, as one may see directly. But it illustrates the point most simply – without some of the technical window dressing – and so the author hopes that the reader will indulge him and pretend that he does not realize this already.) A line bundle on $X^{\prime}$ is then automatically trivial, and a descent datum for $X^{\prime}_{T}/X_{T}$ consists of nothing other than a unit on $(X^{\prime}\times_{X}X^{\prime})_{T}=(\mathbf{A}^{1}_{T})_{K^{\prime}\otimes_{K}K^{\prime}}$ (which is to say, an automorphism of the trivial line bundle) that satisfies the cocycle condition. The problem is that $K^{\prime}\otimes_{K}K^{\prime}$ may well be non-reduced. Indeed, in the situation of main interest for our result – when $K$ is imperfect – this will typically be the case. Therefore a unit on $(\mathbf{A}^{1}_{K^{\prime}\otimes_{K}K^{\prime}})_{T}$ may have arbitrarily high degree in the variable $y$ on $\mathbf{A}^{1}$. Without bounding this degree a priori, we would obtain not representability by a scheme, but by an ind-scheme, because the set of all polynomials is not represented by a scheme. Thus we require some sort of bound on the degree. This will follow from the cocycle condition, which says that our unit $u$ satisfies
$u_{12}u_{23}=u_{13},$ (1.1)
where $u_{ij}$ is the image of $u$ under the “projection” map $K^{\prime}\otimes_{K}K^{\prime}\rightarrow K^{\prime}\otimes_{K}K^{\prime}\otimes_{K}K^{\prime}$ which sends $a\otimes b$ to $a\otimes b$ in the $(i,j)$ coordinates and $1$ in the other coordinate. Looking at the leading terms in (1.1), degree considerations would allow us to conclude that the degree of $u$ is bounded (that $u$ is constant, in fact) if $K^{\prime}\otimes_{K}K^{\prime}\otimes_{K}K^{\prime}$ were an integral domain. Unfortunately, this is not the case (the non-reducedness of this ring being the whole cause of our troubles). But what we will show is a poor man’s version of the domain property, which says that if $a,b\in K^{\prime}\otimes_{K}K^{\prime}$ satisfy $a_{12}b_{23}=0$, then either $a$ or $b$ must vanish. (In fact, we prove something somewhat more general; see Proposition 4.2.) This will then allow us to conclude. The argument in general is similar, but with some additional technical complications.
The structure of this paper is as follows. In §2 we descend the almost-representability of the restricted Picard functor through finite Galois covers. In §3 we discuss the notion of boundedness for a collection of descent data, and show how boundedness of the collection of all possible descent data on certain line bundles implies the almost-representability of the restricted Picard functor. In §4 we use the notion of boundedness to construct the tools required to descend almost-representability of the restricted Picard functor through purely inseparable extension fields of degree $p$. In §5 we prove the almost-representability of the restricted Picard functor for regular $K$-schemes of finite type. Finally, in §6, we study the structure of the restricted Picard functor.
1.1 Notation And Conventions
Throughout this paper, $K$ denotes a field, and $K_{s}$, $K_{\rm{perf}}$, $\overline{K}$ denote separable, perfect, and algebraic closures of $K$, respectively. When it appears, $p>0$ is a prime denoting the characteristic of $K$.
2 Descent through Galois covers
In this section we descend the (almost-)representability of the restricted Picard functor through finite Galois covers of schemes. We first prove the following proposition on the representability of unit groups.
Proposition 2.1.
Let $f\colon X\rightarrow{\rm{Spec}}(K)$ be a normal $K$-scheme of finite type. There is a unique subfunctor ${\rm{Units}}_{X/K}$ of the functor
[…]
with the following two properties:
  • (i)
    ${\rm{Units}}_{X/K}$ is represented by a smooth $K$-scheme.
  • (ii)
    ${\rm{Units}}_{X/K}(T)=(f_{*}{\mathbf{G}}_{\rm{m}})(T)$ for every smooth $K$-scheme $T$.
Furthermore, ${\rm{Units}}_{X/K}$ is a commutative $K$-group scheme such that the $K_{s}$-points of its component group are free of finite rank, and ${\rm{Units}}_{X/K}^{0}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$, where $A\subset\Gamma(X,\mathcal{O}_{X})$ is the ring of elements algebraic over $K$.
Proof.
Let us first prove the following claim. Given a normal $K$-scheme $X$ of finite type, a dense open subscheme $U\subset X$, and an element $a\in\Gamma(U,\mathcal{O}_{U})$ algebraic over $K$, $a$ extends uniquely to a section of $X$. Indeed, let $p\in X\backslash U$ be a codimension $1$ point. Then $a|U$ defines a rational map $g\colon X\dashrightarrow\mathbf{P}^{1}_{K}$ which necessarily extends over $p$. Because $a$ is algebraic over $K$, $g|U$ lands inside a finite closed subscheme $Z$ of $\mathbf{A}^{1}_{K}$, and $g^{-1}(Z)$ therefore defines a closed subscheme of a neighborhood $W$ of $p$ containing $W\cap U$, hence all of $W$. We have thus extended $a$ over a neighborhood of $p$, and the reducedness of $W$ ensures that this extension is unique. Thus we obtain a unique extension of $a$ over each codimension $1$ point of $X$, and because $X$ is normal, this then extends to all of $X$, as claimed.
Now to the assertion of the proposition. The uniqueness of ${\rm{Units}}_{X/K}$ follows from Yoneda’s Lemma, as does the fact that it is a commutative $K$-group scheme. To prove existence and the other assertions, we are free by Galois descent to replace $K$ by a finite Galois extension. First assume that $X$ is separated, and choose a normal compactification $\overline{X}$ of $X$. Let $A:=\Gamma(\overline{X},\mathcal{O}_{\overline{X}})$, a finite $K$-algebra, hence algebraic over $K$. Note that its image in $\Gamma(U,\mathcal{O}_{U})$ consists exactly of the elements algebraic over $K$, thanks to the claim of the first paragraph. We have an inclusion $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})\hookrightarrow f_{*}{\mathbf{G}}_{\rm{m}}$, and we wish to show among other things that this map induces an isomorphism onto ${\rm{Units}}_{X/K}^{0}$. For this we are free to replace $K$ by a finite Galois extension. Extending scalars, we may assume that all of the boundary components $E_{1},\dots,E_{n}$ of $\overline{X}\backslash X$ are geometrically irreducible. For any smooth $K$-scheme $T$, $X_{T}$ is still normal, and a uniformizer of $D_{i}$ is also a uniformizer of $(D_{i})_{T}$. We thus obtain a homomorphism
$\phi\colon(f_{\operatorname{\acute{e}t}})_{*}{\mathbf{G}}_{\rm{m}}\rightarrow\mathbf{Z}^{n}$
given by taking valuations along each of the $D_{i}$, where $f_{\operatorname{\acute{e}t}}$ is the map on étale sites induced by $f$. Then $\ker(\phi)(T)=(\overline{f}_{*}{\mathbf{G}}_{\rm{m}})(T)$ for every smooth $T/K$, where $\overline{f}\colon\overline{X}\rightarrow{\rm{Spec}}(K)$ is the structural morphism. If $A:=\Gamma(\overline{X},\mathcal{O}_{\overline{X}})$, a finite $K$-algebra, then $\overline{f}_{*}{\mathbf{G}}_{\rm{m}}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$. Replacing $K$ by a finite Galois extension, we may assume that $\operatorname{im}(\phi)$ is constant and that every $K$-point in $\operatorname{im}(\phi)$ lifts to a point of $(f_{\operatorname{\acute{e}t}})_{*}{\mathbf{G}}_{\rm{m}}$. So there exist units $u_{1},\dots,u_{n}$ on $X$ whose images freely generate $\operatorname{im}(\phi)$. Then every unit on $X_{T}$ for smooth $T$ is in the free subgroup generated by the $u_{i}$ times a unit extending over $\overline{X}_{T}$, which is to say, an element of $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})(T)$. Thus we obtain a smooth $K$-group scheme ${\rm{Units}}_{X/K}$ which is an extension of a free finite rank constant group scheme by $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$, which computes the units of $X_{T}$ for smooth $T$, and which is a subfunctor of the functor of units (because $u_{i}$ and points of $\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$ define units on $X_{T}$ for any $K$-scheme $T$). This completes the proof when $X$ is separated.
Now consider the general, possibly non-separated case. Let $U\subset X$ be a separated dense open subscheme (for instance, a dense affine open). Replacing $K$ by a finite Galois extension if necessary, we may assume that all of the boundary components of $X\backslash U$ are geometrically irreducible. Let $D_{1},\dots,D_{m}$ be these components. For any smooth scheme $T$, $X_{T}$ is still normal, and a uniformizer of $D_{i}$ is also a uniformizer of $(D_{i})_{T}$. We thus obtain a homomorphism
$\phi\colon{\rm{Units}}_{U/K}\rightarrow\mathbf{Z}^{m}$
given by taking valuations along each of the $D_{i}$, and $\ker(\phi)(T)=(f_{*}{\mathbf{G}}_{\rm{m}})(T)$ for every smooth $T/K$. Because the codomain of $\phi$ is étale, $\ker(\phi)^{0}={\rm{Units}}_{U/K}^{0}$ and $\ker(\phi)$ is smooth. Furthermore, the algebraic subrings of $\Gamma(X,\mathcal{O}_{X})$ and $\Gamma(U,\mathcal{O}_{U})$ agree by the first paragraph. Thus $\ker(\phi)^{0}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$, where $A\subset\Gamma(X,\mathcal{O}_{X})$ is a finite $K$-algebra, hence $\ker(\phi)^{0}$ defines a subfunctor of $f_{*}{\mathbf{G}}_{\rm{m}}$. A similar argument to the one in the first paragraph involving the group freely generated by the $u_{i}$ then shows that $\ker(\phi)$ is a subfunctor of $f_{*}{\mathbf{G}}_{\rm{m}}$, hence ${\rm{Units}}_{X/K}=\ker(\phi)$. That the component group is free of finite rank follows from the analogous assertion for $U$. ∎
In the process of descending Picard functors through Galois covers, we will have to make use of some cohomology of finite groups, and in fact to compute certain functors defined in terms of such groups. We therefore study such functors below.
Definition 2.2.
Let $S$ be a scheme, $\mathscr{M}$ a sheaf of abelian groups on the smooth-étale site of $S$, and $G$ a group acting on $\mathscr{M}$. Then $\mathscr{H}^{i}(G,\mathscr{M})$ is the (étale) sheafification of the presheaf $T\mapsto{\rm{H}}^{i}(G,\mathscr{M}(T))$.
Proposition 2.3.
If $E$ is an étale commutative $K$-group scheme, and $G$ is a finite group acting on $E$, then $\mathscr{H}^{i}(G,E)$ is an étale $K$-group scheme. If $E(K_{s})$ is finitely generated, then $\mathscr{H}^{i}(G,E)(K_{s})$ is finitely generated for $i=0$ and finite for $i>0$.
Proof.
By étale descent, we may assume that $E$ is the constant group scheme associated to a $G$-module $M$. Then $\mathscr{H}^{i}(G,E)$ is the constant $K$-group scheme associated to the group ${\rm{H}}^{i}(G,M)$, which proves the first assertion. For the second, it suffices to note that ${\rm{H}}^{i}(G,M)$ is finitely generated (for instance, by the cocycle definition of group cohomology), and that, for $i>0$, it is also torsion, hence finite. ∎
Proposition 2.4.
For a smooth commutative $K$-group scheme $M$ of finite type equipped with an action by the finite group $G$, $\mathscr{H}^{i}(G,M)$ is represented by a smooth commutative $K$-group scheme of finite type, which is affine if $M$ is so.
Proof.
The functor $\mathcal{Z}^{n}(G,M)$ of $n$-cocycles $G^{n}\rightarrow M$ is readily checked to be representable by the closed subscheme of $M^{G^{n}}$ corresponding to the cocycle condition. We also have coboundary maps $d_{n+1}\colon M^{G^{n}}\rightarrow\mathcal{Z}^{n}(G,M)$. Because $M^{G^{n}}$ is smooth, the quotient group sheaf $Q^{i}:=\mathcal{Z}^{i}(G,M)/{\rm{im}}(d_{i+1})$ on the smooth-étale site of $K$ agrees with the usual fppf quotient group scheme. Replacing $Q^{i}$ by its maximal smooth $K$-subgroup scheme ([CGP, Lem.C.4.1, Rem. C.4.2]) does not change the underlying functor on the smooth-étale site of $K$, so we see that $\mathscr{H}^{i}(G,M)$ is represented by a smooth commutative finite type $K$-group scheme, which is affine if $M$ is. ∎
Proposition 2.5.
Let $M$ be a smooth commutative $K$-group scheme with finitely generated component group, and suppose that $M$ comes equipped with an action by a finite group $G$. Then ${\mathscr{H}}^{i}(G,M)$ is (represented by) a smooth $K$-group scheme with finitely generated component group. For $i>0$, ${\mathscr{H}}^{i}(G,M)$ is finite type, and it is affine if $M^{0}$ is so.
Proof.
Let $E:=M/M^{t}$ be the torsion-free quotient of the component group of $M$. By Galois descent, we may assume that $K$ is separably closed, so that $E$ is constant and the quotient map $M\rightarrow E$ admits a section, so that $M\simeq M^{t}\times E$. We are thus reduced to the étale and finite type cases, which are handled by Propositions 2.3 and 2.4. ∎
Remark 2.6.
One may of course – with appropriate modifications – prove analogues of the above results for cohomology sheaves of group schemes over an arbitrary base. We will never use this, however.
Let us introduce a definition that describes the type of representability that arises in Theorem 1.3.
Definition 2.7.
We say that a sheaf $\mathscr{F}$ of abelian groups on the smooth-étale site of ${\rm{Spec}}(K)$ is almost-representable when it is of the form $G/E$ for a smooth commutative $K$-group scheme $G$ with finitely generated component group and a subsheaf $E\subset G^{t}$ represented by an étale $K$-group scheme $E$ with $E(K_{s})$ free of finite rank.
Lemma 2.8.
Let $E,E^{\prime}$ be étale $K$-group schemes with finitely generated groups of $K_{s}$-points. Then $\operatorname{Ext}^{1}(E,E^{\prime})$ is a torsion group.
Proof.
Filtering $E$ and $E^{\prime}$ between their torsion subgroups and the free quotients, we may assume that $E(K_{s})$ and $E^{\prime}(K_{s})$ are free of finite rank. Then $\operatorname{Ext}^{1}_{K_{s}}(E,E^{\prime})=0$, so $\operatorname{Ext}^{1}(E,E^{\prime})={\rm{H}}^{1}(\operatorname{Gal}(K_{s}/K),\operatorname{Hom}(E(K_{s}),E^{\prime}(K_{s})))$. Since higher Galois cohomology is torsion, the lemma follows. ∎
We will need the following lemma in order to recognize almost-representable sheaves.
Lemma 2.9.
If $G$ is a smooth commutative $K$-group scheme with finitely generated component group, and $E\subset G$ is a subsheaf on the smooth-étale site of ${\rm{Spec}}(K)$ represented by an étale $K$-group scheme with $E(K_{s})$ finitely generated, then $G/E$ is almost-representable.
Proof.
Let $C:=G/G^{t}$ denote the torsion-free quotient of the component group of $G$, and let $\phi\colon E\rightarrow C$ denote the composition $E\subset G\rightarrow C$. First assume that $\phi(E)=0$. Then $E\subset G^{t}$. Let $T\subset E$ denote the torsion subgroup. Then $\overline{G}:=G/T$ is a smooth $K$-group scheme with finitely-generated component group, and $G/E=\overline{G}/(E/T)$, so $G/E$ is almost-representable.
Now assume that $\phi(E)\subset C$ is nonzero. Then $E$ defines an element of $\operatorname{Ext}^{1}(\phi(E),\ker(\phi))$. By Lemma 2.8, this element is killed by some $N>0$. The multiplication by $N$ map induces an isomorphism $\phi(E)\simeq[N]\phi(E)$, so the extension
[…]
splits. Thus there is a $K$-subgroup scheme $H$ of $E$ mapping isomorphically onto $[N]\phi(E)$. Then $G/H$ is an extension of $C/[N]\phi(E)$ by $G^{t}$, hence representable by a scheme $G^{\prime}$ with finitely generated component group. Letting $\overline{E}:=E/H$, we have $G/E\simeq G^{\prime}/\overline{E}$, and $\overline{E}(K_{s})$ has strictly smaller rank than $E(K_{s})$, hence $G^{\prime}/\overline{E}$ is almost-representable by induction. ∎
Lemma 2.10.
Suppose given a short exact sequence of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$
[…]
with $H$ and $Q$ locally finite type group schemes over $K$. Then $\mathscr{G}$ is a smooth $K$-group scheme, and $\pi$ is a smooth morphism.
Proof.
We first show that $\mathscr{G}$ is a scheme. By [Stacks, Prop. 79.10.3], it suffices to show that $\mathscr{G}$ is a separated algebraic space over $K$. First, replacing $H$ and $Q$ by their maximal smooth $K$-subgroup schemes has no effect on the resulting sheaves, so we may assume that both groups are smooth. Choose an étale surjection $f\colon X\rightarrow Q$ such that $f$ lifts to a point $s\in\mathscr{G}(X)$. Then $Y:=\mathscr{G}\times_{Q}X\simeq H\times X$ via $(g,x)\mapsto(gs(x)^{-1},x)$, so this gives a (locally finite type) $K$-scheme admitting an étale surjection onto $\mathscr{G}$. It only remains to show that $\mathscr{G}$ is separated.
We must show that the diagonal map $\Delta\colon\mathscr{G}\hookrightarrow\mathscr{G}\times\mathscr{G}$ is a closed embedding. This may be done after base change to some étale cover of the target. We take as our cover the scheme $Y\times Y$. Thus we wish to verify that the top row of the Cartesian diagram
[…]
is a closed embedding. Now $Y\times_{\mathscr{G}}Y$ is the subscheme of $Y\times Y=(H\times X)^{2}$ defined by the condition $h_{1}s(x_{1})=h_{2}s(x_{2})$. This condition in particular implies that $f(x_{1})=f(x_{2})$, so $Y\times_{\mathscr{G}}Y$ is contained in the preimage $Z\subset Y\times Y$ of the diagonal under the map $Y\times Y\rightarrow Q\times Q$. Because $Q$ – being a locally finite type group scheme over a field – is separated, $Z$ is a closed subscheme of $Y\times Y$. Finally, $Y\times_{\mathscr{G}}Y$ is the preimage of $1\in H(K)$ under the map $Z\rightarrow H$, $(h_{1},x_{1},h_{2},x_{2})\mapsto h_{1}s(x_{1})(h_{2}s(x_{2}))^{-1}$, hence it is a closed subscheme of $Z$.
Replacing a representing object for $\mathscr{G}$ by its maximal smooth $K$-subgroup, we may assume that $\mathscr{G}$ is a smooth $K$-group, and we wish to show that $\pi$ is smooth. This may be checked after an étale base change, hence it is equivalent to check that $Y\simeq H\times X\rightarrow X$ is smooth, and this is clear. ∎
We now establish certain permanence properties for almost-representable sheaves.
Lemma 2.11.
The kernel of a homomorphism between almost-representable sheaves is almost-representable.
Proof.
Let $\phi\colon\mathscr{F}\rightarrow\mathscr{G}$ be the map. Write $\mathscr{F}\simeq F/E_{F}$ and $\mathscr{G}=G/E_{G}$, where $F,G$ are smooth commutative $K$-group schemes with finitely-generated component groups, and $E_{F}\subset F$, $E_{G}\subset G$ are subsheaves represented by étale $K$-group schemes with finitely generated groups of $K_{s}$-points. Let $q_{F}\colon F\rightarrow\mathscr{F}$ denote the quotient map and similarly for $q_{G}\colon G\rightarrow\mathscr{G}$. Let $W$ be the sheaf defined by the following Cartesian diagram:
[…]
The map $q_{G}$ is surjective with kernel isomorphic to $E_{G}$, hence the same holds for $\zeta$. Lemma 2.10 therefore implies that $W$ is a smooth $F$-group scheme. Furthermore, the subsheaf $\eta^{-1}(E_{G})\subset W$ surjects onto $\ker(\phi)$ via $q_{F}\circ\zeta$. Call this surjection $\delta$. Then $\ker(\delta)=\zeta^{-1}(E_{F})$. This is étale with finitely generated component group, because $\ker(\zeta)\simeq\ker(q_{G})=E_{G}$. By Lemma 2.9, therefore, in order to prove the lemma it is enough to show that $U:=\eta^{-1}(E_{G})\subset W$ is a smooth $K$-group scheme with finitely generated component group.
The sheaf $U$ is defined by the following Cartesian diagram:
[…]
It follows that $U$ is computed by the locally finite type $K$-group scheme $S$ defined by the same Cartesian diagram in the category of schemes, and replacing $S$ by its maximal smooth $K$-subgroup scheme, we see that $U$ is a smooth $K$-group scheme. It remains to prove that it has finitely generated component group. Because $U$ is a closed subgroup subscheme of $S$, it is enough to prove the same for $S$.
Because $W$ is an $E_{G}$-torsor over $F$, it has finitely generated component group, hence the same holds for the closed $K$-subgroup scheme $\ker(\eta)_{\rm{gp}}\subset W$, defined as the kernel in the category of $K$-group schemes. (Note: This is not generally the same as the kernel in the category of sheaves on the smooth-étale site of $K$. The relationships is that the latter is the maximal smooth $K$-subgroup scheme of the former.) But $\ker(\eta)_{\rm{gp}}=\ker(S\rightarrow E_{G})_{\rm{gp}}$. It follows that $S$ has finitely generated component group. ∎
Lemma 2.12.
Let $\mathscr{F}$ be a sheaf on the smooth-étale site of $K$, equipped with an action by a finite group $G$. If $\mathscr{F}$ is almost-representable, then so is the sheaf $\mathscr{F}^{G}$ of $G$-invariants.
Proof.
Apply Lemma 2.11 to the map $\mathscr{F}\rightarrow\prod_{g\in G}\mathscr{F}$ defined by $f\mapsto((g-1)f)_{g}$. ∎
Lemma 2.13.
Any subsheaf of an étale sheaf $E$ on the smooth-étale site of $K$ is also étale.
Proof.
[Ros3, Lem. 5.5] says that subsheaves of étale sheaves on the geometrically reduced étale site of ${\rm{Spec}}(K)$ are still étale, and the proof of that lemma also goes through for the smooth-étale site. ∎
Lemma 2.14.
A commutative extension of an almost-representable group sheaf by an affine, finite type group sheaf is almost-representable.
Proof.
Suppose given an extension $\mathscr{F}$ of $G/E$ by $H$ with $H$ an affine $K$-group scheme of finite type, $G$ a smooth commutative $K$-group scheme with finitely generated component group, and $E\subset G$ an étale $K$-group scheme with $E(K_{s})$ free of finite rank. Replacing $H$ by its maximal smooth $K$-subgroup scheme, we may as well assume that $H$ is smooth. Consider the following commutative diagram with exact rows and columns, in which $W$ is defined by the indicated Cartesian diagram:
[…]
By the effectivity of fpppf descent for relatively affine schemes, $W$ is a (smooth) $G$-scheme. Furthermore, the middle row shows that $W$ has finitely generated component group. It follows that $\mathscr{F}\simeq W/E$ is almost-representable. ∎
We may now show that almost-representability of the restricted Picard functor descends through finite Galois covers.
Proposition 2.15.
Let $Y\rightarrow X$ be a finite Galois cover of normal $K$-schemes of finite type. If $\operatorname{Pic}^{+}_{Y/K}$ is almost-representable, then so is $\operatorname{Pic}^{+}_{X/K}$.
Proof.
Let $G$ be the Galois group. Then $Y_{T}\rightarrow X_{T}$ is a $G$-cover for any $K$-scheme $T$, and we have the Hochschild-Serre spectral sequence, functorial in $T$,
$E_{2}^{i,j}={\rm{H}}^{i}(G,{\rm{H}}^{j}(Y_{T},{\mathbf{G}}_{\rm{m}}))\Longrightarrow{\rm{H}}^{i+j}(X_{T},{\mathbf{G}}_{\rm{m}}).$
We thus obtain an exact sequence of sheaves
[…]
Propositions 2.1 and 2.5 imply that ${\rm{H}}^{i}(G,{\rm{Units}}_{Y/K})$ is a smooth affine $K$-scheme for $i>0$. Combining Lemmas 2.12 and 2.11, therefore, we deduce that $\ker(f)$ is almost-representable. Lemma 2.14 then implies that $\operatorname{Pic}^{+}_{X/K}$ is almost-representable. ∎
3 Boundedness
We will ultimately prove the almost-representability of $\operatorname{Pic}^{+}_{X/K}$ by relating line bundles on $X$ to line bundles on a suitable fppf cover $Y$ of $X$ together with descent data. A descent datum for a line bundle amounts to a trivialization of a certain associated line bundle. We will have to understand those trivializations that show up as descent data, and in particular show that they yield a representable functor. The key to accomplishing this is the notion of boundedness for a family of trivializations.
Definition 3.1.
(Boundedness) Let $Z$ be a $K$-scheme, and let $V\subset\Gamma(Z,\mathcal{O}_{Z})$ be a finite-dimensional $K$-vector space. Suppose given a $K$-scheme $T$ and $s\in\Gamma(Z\times_{K}T,\mathcal{O}_{Z\times_{K}T})$. We say that the pair $(T,s)$ is $V$-bounded when $s\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$.
Let $S$ be a $K$-scheme, and let $\mathscr{L},\mathscr{M}$ be line bundles on $Z\times_{K}S$, and ϕ::italic-$\phi\colon\mathscr{L}\xrightarrow{\sim}\mathscr{M}$ an isomorphism. Suppose given a pair $(g,\psi)$ with $g\colon T\rightarrow S$ an $S$-scheme and $\psi\colon({\rm{Id}}_{Z},g)^{*}(\mathscr{L})\xrightarrow{\sim}({\rm{Id}}_{Z},g)^{*}(\mathscr{M})$ an isomorphism. We say that $(g,\psi)$ is $(\phi,V)$-bounded when the pair
$(T,\psi\circ({\rm{Id}}_{Z},g)^{*}(\phi)^{-1})\in\Gamma(Z\times_{K}T,{\mathbf{G}}_{\rm{m}}))$
and its inverse $(T,({\rm{Id}}_{Z},g)^{*}(\phi)\circ\psi^{-1})$ are both $V$-bounded.
Fix $\phi$. We say that a family $F:=\{(g_{\alpha},\psi_{\alpha})\}_{\alpha\in A}$ of pairs as above is bounded when there exists a finite-dimensional $V$ such that every element of the family is $(\phi,V)$-bounded. Note that this is independent of $\phi$ if $S$ is quasi-compact. In particular, we say (without reference to any $\phi$) that $F$ is locally bounded when there is a Zariski cover $\{U_{i}\}_{i\in I}$ of $S$ such that, for each $i$, $U_{i}$ is quasi-compact and $F_{i}:=\{(g_{\alpha,i},\psi_{\alpha,i})\}$ is bounded, where $g_{\alpha,i}:=g_{\alpha}|g_{\alpha}^{-1}(U_{i})$ and $\psi_{\alpha,i}=\psi_{\alpha}|{Z\times_{K}g_{\alpha}^{-1}(U_{i})}$. Note that we require the existence of an isomorphism $\phi$ to make this definition, but the definition is independent of which isomorphism we choose.
Proposition 3.2.
Fix $Z,S,\mathscr{L},\mathscr{M},\phi,V$ as in Definition 3.1 with $S$ a finite type $K$-scheme. Then the set-valued functor on the category of $K$-schemes sending $T$ to the set of all $(\phi,V)$-bounded pairs $(g,\psi)$ with $g\colon T\rightarrow S$ a $K$-morphism and $\psi\colon({\rm{Id}}_{Z},g)^{*}(\mathscr{L})\xrightarrow{\sim}({\rm{Id}}_{Z},g)^{*}(\mathscr{M})$ an isomorphism, is represented by a finite type $K$-scheme ${\rm{Isom}}_{V}$. Furthermore, for $V\subset W\subset\Gamma(Z,\mathcal{O}_{Z})$ finite-dimensional $K$-subspaces, the natural inclusion ${\rm{Isom}}_{V}\subset{\rm{Isom}}_{W}$ is a closed immersion.
Proof.
We first prove the representability. Once one has specified $g$, to give a $(\phi,V)$-bounded isomorphism $\psi$ is the same as giving a unit $u\in\Gamma(Z\times_{K}T,{\mathbf{G}}_{\rm{m}})$ such that $u,u^{-1}\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$. The datum of an element $f\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$ is the same as a $K$-morphism from $T$ to the vector group $\underline{V}$ defined by $V$. Thus to specify $\psi$ is the same as specifying a pair $(f,f^{\prime})$ of $K$-morphisms $T\rightarrow\underline{V}$ such that $ff^{\prime}=1$. We claim that this condition defines a closed subscheme $Y\subset\underline{V}\times\underline{V}$, in which case we see that ${\rm{Isom}}_{V}=S\times Y$ is a scheme.
To see that $ff^{\prime}=1$ indeed defines a closed subscheme of $\underline{V}\times\underline{V}$, choose a basis :=$\mathscr{B}:=\{v_{1},\dots,v_{n}\}$ of $V$, with $1\in\mathscr{B}$ if $1\in V$. Let $V^{\prime}\subset\Gamma(Z,\mathcal{O}_{Z})$ be a finite-dimensional subspace with $V\cdot V\subset V^{\prime}$ and $1\in V^{\prime}$. Choose a basis :=$\mathscr{B}^{\prime}:=\{1,b_{2},\dots,b_{m}\}$ of $V^{\prime}$. For each $1\leq i,j\leq n$, write
$v_{i}v_{j}=\sum_{1\leq k\leq m}\lambda_{ijk}b_{k}$ (3.1)
for unique $\lambda_{ijk}\in K$. Then $\underline{V}\simeq\mathbf{A}^{n}$ via the coefficients of the $v_{i}$, and if we let $d_{i},e_{i}$ denote the coordinates on $\mathbf{A}^{n}$, then the condition $ff^{\prime}=1$ is the same as expanding the left side in terms of the basis $\mathscr{B}^{\prime}$ using the equations (3.1), and then equating the coefficients of the $b_{k}$ on both sides. This yields a collection of equations defining a closed subscheme of $\underline{V}^{2}$. This completes the proof that ${\rm{Isom}}_{V}$ is representable.
To prove that ${\rm{Isom}}_{V}\hookrightarrow{\rm{Isom}}_{W}$ is a closed immersion, we note that it suffices to show that the map $\underline{V}^{2}\hookrightarrow\underline{W}^{2}$ is a closed immersion, and this is well-known (and easy): Just complete the chosen basis $\mathscr{B}$ for $V$ to a basis $\mathscr{B}\sqcup\mathscr{C}$ for $W$, and then $\underline{V}$ is defined by the conditions that the coefficients of the elements of $\mathscr{C}$ vanish. ∎
Because we will analyze line bundles on a scheme $X$ by considering descent data on some fppf cover $f\colon Y\rightarrow X$, the following definition will play an important role in proving the representability of $\operatorname{Pic}^{+}_{X/K}$.
Definition 3.3.
Let $f\colon Y\rightarrow X$ be a morphism of $K$-schemes. We define the presheaf of abelian groups ${\rm{DesCan}}_{f}^{\rm{naive}}$ on the smooth-étale site of $K$ by the formula
$T\mapsto\ker\left((\pi_{1})_{T}^{*}-(\pi_{2})_{T}^{*}\colon\operatorname{Pic}(Y\times_{K}T)\rightarrow\operatorname{Pic}(Y\times_{X}Y\times_{K}T)\right),$
where $\pi_{i}\colon Y\times_{X}Y\rightarrow Y$ are the projections. We define ${\rm{DesCan}}_{f}$ to be the (étale) sheafification of ${\rm{DesCan}}_{f}^{\rm{naive}}$. We also denote these functors by ${\rm{DesCan}}_{Y/X}^{\rm{naive}}$ and ${\rm{DesCan}}_{Y/X}$, respectively.
The importance of boundedness arises from the following proposition.
Proposition 3.4.
Suppose given an fpqc morphism of connected normal $K$-schemes $f\colon Y\rightarrow X$. Assume that ${\rm{DesCan}}_{Y/X}$ is almost-representable, and choose an isomorphism ${\rm{DesCan}}_{Y/X}\xrightarrow{\sim}G/E$ exhibiting it as such. Choose an étale cover $U\rightarrow G$ such that the element of ${\rm{DesCan}}_{Y/X}(G)$ corresponding to the quotient map $G\rightarrow G/E$ is represented on $U$ by a line bundle
$\mathscr{L}\in\ker(\pi_{1}^{*}-\pi_{2}^{*}\colon\operatorname{Pic}(Y\times_{K}U)\rightarrow\operatorname{Pic}(Y\times_{X}Y\times_{K}U)).$
Let $F$ be the family of pairs $(u,\psi)$ where $u\in U(K_{s})$ and $\psi\colon(\pi_{1})_{K_{s}}^{*}(u^{*}\mathscr{L})\xrightarrow{\sim}(\pi_{2})_{K_{s}}^{*}(u^{*}\mathscr{L})$ is an isomorphism satisfying the cocycle condition
$\pi_{12}^{*}(\psi)\circ\pi_{23}^{*}(\psi)=\pi_{13}^{*}(\psi),$
where $\pi_{ij}\colon Y\times_{X}Y\times_{X}Y\rightarrow Y\times_{X}Y$ is projection onto the $(i,j)$ factors. If $F$ is locally bounded, then $\operatorname{Pic}^{+}_{X/K}$ is almost-representable.
Proof.
Let $P$ be the presheaf on the smooth-étale site of ${\rm{Spec}}(K)$ defined by the following Cartesian diagram:
[…]
so that the (étale) sheafification $\mathscr{F}$ of $P$ lives in the following Cartesian diagram:
[…]
We will first show that there is a locally finite type, geometrically reduced $K$-algebraic space computing $\mathscr{F}$. In order to prove this, we choose an open cover $\{U_{i}\}_{i\in I}$ of $U$ by finite type $K$-schemes . Then, for each $i$, the family of $(u,\psi)$ as in the proposition such that $u\in U_{i}(K_{s})$ is bounded.
Let $P_{i}$ denote the preimage of $U_{i}$ in $P$. For a smooth $K$-scheme $T$, to give a $T$-point of $P_{i}$ is the same as giving a morphism $g\colon T\rightarrow U_{i}$ together with a line bundle (up to isomorphism) $\mathscr{M}$ on $X\times_{K}T$ such that $({\rm{Id}}_{Y},g)^{*}(\mathscr{L})=f_{T}^{*}(\mathscr{M})$. By the theory of faithfully flat descent, this is (functorially) the same as the information of the morphism $g$ together with a descent datum (up to isomorphism) on $({\rm{Id}}_{Y},g)^{*}(\mathscr{L})$ with respect to the map $f_{T}$. Fix an isomorphism $\phi\colon(\pi_{1})_{U}^{*}(\mathscr{L})\xrightarrow{\sim}(\pi_{2})_{U}^{*}(\mathscr{L})$. Then a descent datum over $T$ is described by a unit $u\in\Gamma(Y\times_{X}Y\times_{K}T,{\mathbf{G}}_{\rm{m}})$ such that $u\cdot g^{*}(\phi)$ satisfies the cocycle condition. By assumption, the set of such $u$, and the set of $u^{-1}$ for such $u$, is bounded when we restrict to the case $T={\rm{Spec}}(K_{s})$ – that is, it is contained in $V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$ for some finite-dimensional $V\subset\Gamma(Y\times_{X}Y,\mathcal{O}_{Y\times_{X}Y})$. For a general $T$, locally on $T$ one has that $u,u^{-1}\in W\subset\Gamma(Y\times_{X}Y,\mathcal{O}_{Y\times_{X}Y})$ for some finite-dimensional $W\supset V$ (which $W$ is allowed to vary with $T$ and with the open subset of $T$ over which one is working locally). Thus the descent datum for $({\rm{Id}}_{Y},g)^{*}(\mathscr{L})$ lies in ${\rm{Isom}}_{W}$ in the notation of Proposition 3.2, while the restriction to each $K_{s}$-point of $T$ factors through the closed subscheme ${\rm{Isom}}_{V}$. Because $T$ is smooth, $T(K_{s})$ is dense in $T$, so it follows that all of $T$ maps into ${\rm{Isom}}_{V}$. So we see that every descent datum for $\mathscr{L}$ arises from a morphism $T\rightarrow{\rm{Isom}}_{V}\times U$, and we claim that the condition for the descent datum to define a cocycle defines a closed subscheme of ${\rm{Isom}}_{V}\times U$. Indeed, the cocycle condition on $u\cdot g^{*}(\phi)$ is described by the condition
$\pi_{12}^{*}(u)\pi_{23}^{*}(u)=g^{*}(C)\pi_{13}^{*}(u),$ (3.2)
where $C:=\pi_{13}^{*}(\phi)\pi_{23}^{*}(\phi)^{-1}\pi_{12}^{*}(\phi^{-1})$. Choose a finite-dimensional subspace $H\subset\Gamma(Y\times_{X}Y\times_{X}Y,\mathcal{O})$ containing $C\cdot\pi_{13}^{*}(V)$ and $\pi_{12}^{*}(V)\cdot\pi_{23}^{*}(V)$, and choose a basis $\mathscr{B}$ for $H$. Then (3.2) reduces to the equations equating the coefficients of each element of $\mathscr{B}$ on both sides of (3.2). This condition defines a closed subscheme of ${\rm{Isom}}_{V}\times U$, as claimed.
We have thus obtained for each finite-dimensional $W\subset\Gamma(Y\times_{X}Y,\mathcal{O})$ a finite type $U$-scheme $r\colon S^{\prime}_{W}\rightarrow U$ which is a subfunctor of the functor of all descent data for $\mathscr{L}|Y\times_{K}U$, in such a manner that the functor of all descent data is the filtered direct limit of the $S^{\prime}_{W}$, and such that, for each $i$, there is a finite-dimensional $V_{i}$ such that, whenever $V_{i}\subset W$, $S^{\prime}_{W,i}:=S^{\prime}_{W}|r^{-1}(U_{i})$ defines all descent data on $\mathscr{L}|Y\times_{K}U_{i}$ over any smooth $K$-scheme $T$. Replacing $S^{\prime}_{W,i}$ by its maximal geometrically reduced closed subscheme $S_{W,i}$ (see [CGP, Lem.C.4.1]) does not alter its $T$-points for smooth $T$. Since $S^{\prime}_{W_{1},i}\subset S^{\prime}_{W_{2},i}$ is a closed subscheme with the same $K_{s}$-points when $V_{i}\subset W_{1}\subset W_{2}$, we see that $S_{W,i}$ is independent of $W\subset V_{i}$. We denote this common geometrically reduced finite type $U_{i}$-scheme by $S_{i}$. The $S_{i}$ together to yield a finite type $U$-scheme $S$ that is geometrically reduced over $K$ and which computes the functor of descent data for $\mathscr{L}$ from $Y$ to $X$ over any smooth $K$-scheme $T$.
[…]
We claim that $H:={\rm{Units}}_{Y/K}/{\rm{Units}}_{X/K}$ is a smooth $K$-group scheme. Indeed, to verify this, it suffices to prove that the map on component groups induced by the inclusion
${\rm{Units}}_{X/K}\hookrightarrow{\rm{Units}}_{Y/K}$
is injective. For this, we may pass to a finite Galois extension and thereby suppose that we have a unit $u$ on $X$ that does not live in ${\rm{Units}}_{X/K}^{0}$, and we must show that its pullback to $Y$ does not live in ${\rm{Units}}_{Y/K}^{0}$. By Proposition 2.1, $u$ is in the identity component of $X$ precisely when it is algebraic over $K$, and similarly for units on $Y$. Thus what we must show is that, if $u$ is not algebraic, then neither is its pullback to $Y$. But this follows from the fact that the pullback map $f^{*}\colon\Gamma(X,\mathcal{O}_{X})\rightarrow\Gamma(Y,\mathcal{O}_{Y})$ is injective, because $f$ is faithfully flat.
It follows that $\mathscr{F}$ is computed by the geometrically reduced finite type $U$-algebraic space $\mathscr{S}:=S/H$ with structure map $r^{\prime}\colon\mathscr{S}\rightarrow U$. Furthermore, $\mathscr{F}_{i}$ is computed by the finite type $U_{i}$-algebraic space $\mathscr{S}^{\prime}_{W,i}:=S^{\prime}_{W,i}/H$ for any finite-dimensional $V\subset W_{i}$. Let $R\subset U\times U$ be the étale equivalence relation such that $U/R=G$. Because $G$ is a locally finite type $K$-group scheme, it is separated, so $R\subset U\times U$ is a closed subscheme. Then $\iota\colon R^{\prime}:=(r^{\prime}\times r^{\prime})^{-1}(R)\hookrightarrow\mathscr{S}\times\mathscr{S}$ is an et́ale equivalence relation and a closed embedding, and we let $\mathscr{X}:=\mathscr{S}/R^{\prime}$, a locally finite type, separated (because $\iota$ is a closed embedding), geometrically reduced $K$-algebraic space that computes the sheaf $\mathscr{G}$ on the smooth-étale site of ${\rm{Spec}}(K)$ which sits in the following Cartesian diagram
[…]
Note also that, because $\mathscr{S}$ is of finite type over $U$, $\mathscr{X}$ is of finite type over $G$. Similarly, we define for each finite-dimensional $W$ a quotient $\mathscr{X}^{\prime}_{W,i}$ of $\mathscr{S}^{\prime}_{W,i}$ that is an algebraic space whose maximal geometrically reduced closed sub-algebraic space is $\mathscr{X}|b^{-1}(a(U_{i}))$.
Let $D_{i}:=a(U_{i})\subset{\rm{DesCan}}_{Y/X}$, an open subset. The group structure on $U/R=G$, together with that on (isomorphism classes of) line bundles on $X$ descending $\mathscr{L}$, yields a multiplication map for any $i,j,k$ and $W_{1},W_{2}$
$\mathscr{X}^{\prime}_{W_{1},i}\times\mathscr{X}^{\prime}_{W_{2},j}\rightarrow\mathscr{X}^{\prime}_{W_{3},{\bf k}},$
where ${\bf k}\subset I$ is a finite subset such that the union of the $U_{k}$ contains $D_{i}\cdot D_{j}$, and $W_{3}$ is sufficiently large depending on $W_{1},W_{2}$ (and $\mathscr{X}_{W_{3},{\bf k}}$ has the obvious meaning). In this manner one obtains a group structure on the colimit (over all $W,i$) of the $\mathscr{X}^{\prime}_{W,i}$, hence on the colimit of the maximal geometrically reduced closed subschemes, which is $\mathscr{X}$. Thus $\mathscr{X}$ is a geometrically reduced, separated, locally finite type $K$-group algebraic space, so it is a smooth commutative $K$-group scheme [Stacks, Prop. 79.10.3]. Because the map $m\colon\mathscr{X}\rightarrow G$ is of finite type, $\mathscr{X}$ has finitely generated component group. Thus, $\operatorname{Pic}^{+}_{X/K}\simeq\mathscr{X}/E$ is almost-representable. ∎
Lemma 3.5.
The intersection of the functor of units on $X$ with ${\rm{Units}}_{Y/K}$ is ${\rm{Units}}_{X/K}$.
Proof.
We are free to extend scalars and thereby assume that the component group $E$ of ${\rm{Units}}_{Y/K}$ is constant and that every component arises from a unit on $Y$ (that is, the map ${\rm{Units}}_{Y/K}(K)\rightarrow E(K)$ is surjective), and similarly for $X$. Let $B\subset\Gamma(Y,\mathcal{O}_{Y})$ be the subring of elements algebraic over $K$, so ${\rm{Units}}^{0}_{Y/K}=\operatorname{R}_{B/K}({\mathbf{G}}_{\rm{m}})$, and similarly let $A$ be the ring of elements of $\Gamma(X,\mathcal{O}_{X})$ algebraic over $K$, so ${\rm{Units}}^{0}_{X/K}=\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})$. Let $T$ be a $K$-scheme, and suppose given a unit $u\in{\rm{Units}}_{Y/K}(T)$ such that $u$ is the pullback of a unit on $X_{T}$. We must show that $u$ comes from an element of ${\rm{Units}}_{X/K}(T)$.
Working locally, we may assume that $T={\rm{Spec}}(R)$ and that $u$ lies in a single component $C$ of ${\rm{Units}}_{Y/K}$. Choose a unit $v$ on $Y$ defining a point in $C(K)$. Then $C(R)=u(R\otimes_{K}B)^{\times}$, so $v$ lies in the $R$-module $u(R\otimes_{K}B)$. By descent theory, an element of $u(R\otimes_{K}B)$ is the pullback of a section of $X_{R}$ precisely when its pullbacks along the two projections $(Y\times_{X}Y)_{R}\rightarrow Y_{R}$ agree. Because $R$ is $K$-flat, the elements of $u(R\otimes_{K}B)$ whose pullbacks agree are $R\otimes_{K}M$, where $M\subset uB$ are the elements of $uB$ whose pullbacks agree – that is, that descend to sections of $\Gamma(X,\mathcal{O}_{X})$.
If $M=0$, then $v=0$, a contradiction, so $M\neq 0$. Let $0\neq m\in M$. Then $f^{*}(m)\in uB\backslash\{0\}$. Because $Y$ is connected and reduced, $B$ is a field, so $f^{*}(m)$ is a unit of $Y$, hence $m$ is a unit of $X$. Replacing $u$ by $f^{*}(m)$, therefore, we may assume that $u$ descends to a unit of $X$, hence lives in ${\rm{Units}}_{X/K}$. So modifying by $u$, we have reduced to the case in which $C={\rm{Units}}_{Y/K}^{0}=\operatorname{R}_{B/K}({\mathbf{G}}_{\rm{m}})$. Then $M=A$, so an element of $(R\otimes_{K}B)^{\times}$ which is pulled back from $X$ lies in $(R\otimes_{K}A)^{\times}={\rm{Units}}_{X/K}^{0}(R)$. ∎
4 Descent through purely inseparable extensions
In this section we will prove the results required in order to descend the almost-representability of the restricted Picard functor through purely inseparable extensions (although we postpone the proof that representability descends through such extensions until the next section). This may be regarded as the technical heart of the entire paper. We will proceed by verifying the hypotheses of Proposition 3.4. We begin by verifying the almost-representability of ${\rm{DesCan}}$ in this setting.
Proposition 4.1.
Let $X$ be an affine $K$-scheme, $L/K$ a finite purely inseparable extension field, and $Y:=(X_{L})_{\rm{red}}$. Then ${\rm{DesCan}}_{Y/X}\simeq\operatorname{R}_{L/K}(\operatorname{Pic}^{+}_{Y/L})$.
Proof.
Because $L/K$ is radicial, $\operatorname{R}_{L/K}$ induces an equivalence of étale sites on $T$ and $T_{L}$ for any $K$-scheme $T$. It follows that $\operatorname{R}_{L/K}$ commutes with étale sheafification, so it suffices to show that one has an isomorphism
${\rm{DesCan}}_{Y/X}^{\rm{naive}}\simeq\operatorname{R}_{L/K}((\operatorname{Pic}^{+}_{Y/L})^{\rm{naive}})$
when one restricts both sides to the category of affine smooth $K$-schemes. This amounts to showing that, for any smooth affine $T/K$, and every $\mathscr{L}\in\operatorname{Pic}(Y_{T})$, one has $\pi_{1}^{*}(\mathscr{L})\simeq\pi_{2}^{*}(\mathscr{L})$, where $\pi_{i}$ are the base changes to $T$ of the projection maps $Y\times_{X}Y\rightarrow Y$. By [Ros1, Lem. 2.2.9] applied with $G={\mathbf{G}}_{\rm{m}}$ and $i=1$, the embedding $((Y\times_{X}Y)_{T})_{\rm{red}}\hookrightarrow(Y\times_{X}Y)_{T}$ induces an isomorphism on Picard groups. It therefore suffices to prove that the two maps
[…]
agree. Because $L/K$ is purely inseparable, the difference between the two projections $L\rightarrow L\otimes_{K}L$ is nilpotent, hence the two projections
${\rm{Spec}}(L\otimes_{K}L)_{\rm{red}}\rightarrow{\rm{Spec}}(L)$
agree. Because $Y$ is a closed subscheme of $X_{L}$, it follows that the same holds for the projections in (4.1). ∎
It remains to verify the local boundedness of the collection of descent data associated to a finite purely inseparable extension $L/K$. The following lemma – akin to an integral domain property for tensor powers of rings – is simple, but lies at the heart of the proof of Theorem 1.3, and in particular, of the boundedness of the collection of descent data for line bundles on a scheme required in Proposition 3.4 in order to descend representability.
Proposition 4.2.
Let $B$ be a finite free $A$-algebra, and let $\pi_{12},\pi_{23}\colon B\otimes_{A}B\rightarrow B\otimes_{A}B\otimes_{A}B$ denote the maps $b\otimes b^{\prime}\mapsto b\otimes b^{\prime}\otimes 1$, $1\otimes b\otimes b^{\prime}$, respectively. Let $I,I^{\prime},I^{\prime\prime}$ be ideals of $A$ such that, for $b,b^{\prime}\in B$, if $bb^{\prime}\in I^{\prime\prime}B$, then either $b\in IB$ or $b^{\prime}\in I^{\prime}B$. Then for $\alpha,\beta\in B\otimes_{A}B$, if $\pi_{12}(\alpha)\pi_{23}(\beta)\in I^{\prime\prime}(B\otimes_{A}B\otimes_{A}B)$, then either $\alpha\in I(B\otimes_{A}B)$ or $\beta\in I^{\prime}(B\otimes_{A}B)$.
Proof.
Write $B=\oplus_{i=1}^{n}Ab_{i}$ for some $b_{1},\dots,b_{n}\in B$, so
$B\otimes_{A}B\otimes_{A}B=\bigoplus_{i=1}^{n}(B\otimes_{A}B\otimes b_{i}).$
Write $\beta=\sum_{i=1}^{n}(\gamma_{i}\otimes b_{i})$ with $\gamma_{i}\in B$. We then have $\pi_{23}(\beta)=\sum_{i=1}^{n}(1\otimes\gamma_{i}\otimes b_{i})$, so $\pi_{12}(\alpha)\pi_{23}(\beta)=\sum_{i=1}^{n}[\alpha(1\otimes\gamma_{i})\otimes b_{i}]$. Because $B\otimes_{A}B\otimes_{A}B$ is a free left $B\otimes_{A}B$-module over the $1\otimes 1\otimes b_{i}$, to prove the lemma we must show that, if $\alpha\in B\otimes B$ is not in the ideal generated by $I$ and $\gamma\in B$ is not in $I^{\prime}B$, then $\alpha(1\otimes\gamma)\notin I^{\prime\prime}(B\otimes_{A}B)$. Because multiplication by $1\otimes B$ makes $B\otimes_{A}B$ into a free $B$-module, this is the same as asserting that, for $b,b^{\prime}\in B$, $bb^{\prime}\in I^{\prime\prime}B\Longrightarrow b\in IB$ or $b^{\prime}\in I^{\prime}B$, and this is true by assumption. ∎
We shall also require the following lemma about flatness of reductions of base changes to larger fields.
Lemma 4.3.
[…]
Proof.
We may assume that $X={\rm{Spec}}(R)$ with $R$ a normal domain. Let $F:={\rm{Frac}}(R)$, and view $F$ and $L$ as lying in a fixed algebraic closure of $F$. We will prove:
  • (1)
    If $L\subset F$, then the composition $R\rightarrow R_{L}\rightarrow(R_{L})_{\rm{red}}$ is an isomorphism.
  • (2)
    If $L\cap F=K$, then $R_{L}$ is reduced.
These two assertions suffice to establish the lemma, as - because $[L:K]=p$ we must be in either situation (1) or (2).
First let us prove (1). The map $R\rightarrow(R_{L})_{\rm{red}}$ is injective because $R$ injects into $R_{L}$. The key point is therefore surjectivity. We first note that surjectivity holds in the case $R=L$. Indeed, $(L\otimes_{K}L)_{\rm{red}}$ is a finite, reduced, radicial extension of $L$, hence a finite extension field $E/L$. But the map $L\rightarrow L\otimes_{K}L$ admits a section, namely, the multiplication map, hence $L\rightarrow E$ also admits a section, so $E=L$. For general $R$, because $R$ is a normal $K$-algebra and $L\subset F$ is algebraic over $K$, one has $L\subset R$. To prove surjectivity of $R\rightarrow(R\otimes_{K}L)_{\rm{red}}$, we only need show that anything of the form $1\otimes\ell$ with $\ell\in L$ lies in the image, and in fact any such element lies in the image of $L\subset R$ because of the already-treated case $R=L$. This proves (1).
[…]
The following somewhat technical lemma will be used in conjunction with Proposition 4.2 to prove the boundedness of the descent data arising from descent through a finite purely inseparable extension.
Lemma 4.4.
Let $A$ be a Japanese DVR with maximal ideal $\mathfrak{m}$, and let $B$ be a finite free local $A$-algebra which is an integral domain. Assume that $B^{m}:=\{b^{m}\mid b\in B\}\subset A$ for some $m>0$. Then there exists $r>0$ such that, for $b,b^{\prime}\in B$, if $b,b^{\prime}\notin\mathfrak{m}B$, then $bb^{\prime}\notin\mathfrak{m}^{r}B$.
Proof.
Let $\widetilde{B}$ be the normalization of $B$, which is also the normalization of $A$ inside ${\rm{Frac}}(B)$. Because $A$ is Japanese, $\widetilde{B}$ is a finite $A$-module, hence also finite over $B$. It follows that there exists $0\neq a\in B$ such that $a\widetilde{B}\subset B$. Replacing $a$ by $a^{m}$, we may assume that $a\in A$. We claim that $\widetilde{B}$ is a DVR. Indeed, it is a Dedekind domain. Suppose that it admits two distinct maximal ideals $\mathfrak{n}_{1},\mathfrak{n}_{2}$, and let $x\in\mathfrak{n}_{1}-\mathfrak{n}_{2}$. Replacing $x$ by $x^{m}$, we may assume that $x\in A$. But each maximal ideal of $\widetilde{B}$ lies over $\mathfrak{m}$, so $x$ cannot lie in one maximal ideal but not another. Thus $\widetilde{B}$ admits a unique maximal ideal, hence is a DVR.
Let $v:={\rm{ord}}_{\mathfrak{m}}(a)$. We claim that, for $b\in B$, if $b\notin\mathfrak{m}B$, then $b\notin\mathfrak{m}^{v+1}\widetilde{B}$. To see this, let $\pi\in A$ be a uniformizer. If $b\in\mathfrak{m}^{v+1}\widetilde{B}$, then $b/\pi^{v+1}\in\widetilde{B}$, hence $ab/\pi^{v+1}\in B$, or $ab\in\pi^{v+1}B$. Because $B$ is a free $A$-module, this implies that $b\in\pi B$, proving the claim.
Now we claim that we may take $r=2(v+1)$ in the lemma. Indeed, if $b,b^{\prime}\in B$ satisfy $bb^{\prime}\in\mathfrak{m}^{2(v+1)}B\subset\mathfrak{m}^{2(v+1)}\widetilde{B}$, then – using the fact that $A\rightarrow\widetilde{B}$ is a local homomorphism of DVR’s – we conclude that either $b$ or $b^{\prime}$ lies in $\mathfrak{m}^{v+1}\widetilde{B}$; say $b$ does without loss of generality. Then by the claim of the previous paragraph, we deduce that $b\in\mathfrak{m}B$. ∎
We may now verify the local boundedness of the descent data for a purely inseparable extension of degree $p$. For the notion of local boundedness of a family of isomorphisms, see Definition 3.1.
Proposition 4.5.
Let $L/K$ be a purely inseparable, degree $p$ extension of fields, $X$ a normal separated $K$-scheme of finite type, and $Y:=(X_{L})_{\rm{red}}$. Let $T$ be a smooth $K$-scheme, and $\mathscr{L}$ a line bundle on $Y_{T}$ such that $\pi_{1}^{*}(\mathscr{L})\simeq\pi_{2}^{*}(\mathscr{L})\in\operatorname{Pic}((Y\times_{X}Y)_{T})$. Then the collection of all descent data for $\mathscr{L}_{S}$ with respect to the fppf cover $Y_{S}\rightarrow X_{S}$ is locally bounded as $S$ varies over all $T$-schemes smooth over $K$.
Proof.
We may extend scalars and assume that $K$ is separably closed. Further, we may work locally on $T$ and thereby assume that $T$ is quasi-compact, and in this case we must prove boundedness of the collection of descent data. Because the $K$-points of $S$ are Zariski dense in $S$ for any $K$-smooth $S$, we may assume that $S={\rm{Spec}}(K)$. To prove boundedness of the descent data, recall that a descent datum is an isomorphism between the pullbacks of $\mathscr{L}$ along the two projections $Y\times_{X}Y\rightrightarrows Y$ which additionally satisfies the cocycle condition. Fix an isomorphism $\phi\colon\pi_{1}^{*}(\mathscr{L})\xrightarrow{\sim}\pi_{2}^{*}(\mathscr{L})$. Let $s\colon{\rm{Spec}}(K)\rightarrow T$ be the $T$-scheme structure on $S={\rm{Spec}}(K)$. Let $\pi_{ij}\colon Y\times_{X}Y\times_{X}Y\rightarrow Y\times_{X}Y$ denote the various projections. Then the cocycle condition is equivalent to the equation
$\pi_{12}^{*}(u)\pi_{23}^{*}(u)=C\pi_{13}^{*}(u),$
where $C:=s^{*}(C^{\prime})$ with $C^{\prime}:=\pi_{13}^{*}(\phi)\pi_{23}^{*}(\phi)^{-1}\pi_{12}^{*}(\phi)^{-1}\in\Gamma((Y\times_{X}Y\times_{X}Y)\times_{K}T,{\mathbf{G}}_{\rm{m}})$. Because $T$ is quasi-compact, there is a finite-dimensional $W\subset\Gamma((Y^{\times_{X}3}),\mathcal{O})$ such that $C^{\prime},(C^{\prime})^{-1}\in W_{T}$. Now we forget the some of the particulars of our underlying situation, and we simply prove the following: For any finite-dimensional $W\subset\Gamma(Y^{\times_{X}3},\mathcal{O})$, there is a finite-dimensional $V\subset\Gamma(Y\times_{X}Y,\mathcal{O})$ such that any $u\in\Gamma(Y\times_{X}Y,\mathcal{O})$ satisfying
$\pi_{12}^{*}(u)\pi_{23}^{*}(u)\in\pi_{13}^{*}(u)W$ (4.2)
lies in $V$. Then we see that all of the $(t,u)$ defining a cocycle satisfy $u,u^{-1}\in V$.
To prove the claim, we first choose a normal compactification $\overline{X}$ of $X$, and let $\overline{Y}:=(\overline{X}_{L})_{\rm{red}}$, which is a compactification of $Y$ living in a commutative diagram
[…]
with $\overline{f}$ radicial and flat by Lemma 4.3. Let $D_{1},\dots,D_{n}$ be the boundary divisors of $\overline{X}$ (that is, those divisors not contained in $X$), and let $D:=\sum_{i=1}^{n}D_{i}$. For each $N>0$, consider the coherent sheaf $\mathcal{O}_{\overline{X}}(ND)$. Let $g\colon\overline{Y}\times_{\overline{X}}\overline{Y}\rightarrow\overline{X}$ be the natural map, and consider the open embedding $j\colon Y\times_{X}Y\hookrightarrow\overline{Y}\times_{\overline{X}}\overline{Y}$. Because the support of $g^{-1}(ND)$ is disjoint from $Y\times_{X}Y$, one has a canonical isomorphism $j^{*}g^{*}(\mathcal{O}_{\overline{X}}(ND))\xrightarrow{\sim}\mathcal{O}_{Y\times_{X}Y}$ which is adjoint to a map $g^{*}(\mathcal{O}_{\overline{X}}(ND))\rightarrow j_{*}(\mathcal{O}_{Y\times_{X}Y})$. Let $\mathscr{G}_{N}$ denote the image of this map, a coherent sheaf on $\overline{Y}\times_{\overline{X}}\overline{Y}$. We will show that, for some $N>0$ depending on $W$, every $u$ satisfying (4.2) lies in ${\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},\mathscr{G}_{N})\subset{\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},j_{*}\mathcal{O}_{Y\times_{X}Y})={\rm{H}}^{0}(Y\times_{X}Y,\mathcal{O}_{Y\times_{X}Y})$. Because $\overline{Y}\times_{\overline{X}}\overline{Y}$ is proper over $K$ and $\mathscr{G}_{N}$ is coherent, this will show that the set of such $u$ lies in a finite-dimensional $K$-vector space.
To begin with, we claim that being in $\mathscr{G}_{N}$ is determined by codimension-one points. More precisely, let $Z=\coprod{\rm{Spec}}(\mathcal{O}_{s})$ be the disjoint union of the (spectra of the) local rings of all codimension-one points of $\overline{Y}\times_{\overline{X}}\overline{Y}$, and let $h\colon Z\hookrightarrow\overline{Y}\times_{\overline{X}}\overline{Y}$ be the natural map. Then we claim that a section of ($j_{*}$ of) the structure sheaf of $Y\times_{X}Y$ lies in $\mathscr{G}_{N}$ if and only if the same holds upon pulling back to $Z$. This will follow if we can show that the adjunction map $g^{*}(\mathcal{O}(ND))\rightarrow h_{*}h^{*}(g^{*}(\mathcal{O}(ND)))$ is an isomorphism, and similarly with $g^{*}(\mathcal{O}(ND))$ replaced by $\mathcal{O}_{\overline{Y}\times_{\overline{X}}\overline{Y}}$. To check this, note that the analogous statement holds for $\mathcal{O}_{\overline{X}}(ND)$ and the structure sheaf on $\overline{X}$ because $\overline{X}$ is normal: If we define $R$ to be the disjoint union of the spectra of the local rings of $\overline{X}$ at codimension-one points, and $r\colon R\hookrightarrow\overline{X}$ the inclusion, then the corresponding adjunction map is an isomorphism for $\mathcal{O}_{\overline{X}}(ND)$ and for $\mathcal{O}_{\overline{X}}$. Now we claim that if a quasi-coherent sheaf $\mathscr{F}$ on $\overline{X}$ is such that $\mathscr{F}\rightarrow r_{*}r^{*}(\mathscr{F})$ is an isomorphism, then the same holds for the map $g^{*}(\mathscr{F})\rightarrow h_{*}h^{*}(g^{*}(\mathscr{F}))$. To see this, consider the following commutative diagram, which is even Cartesian because $g$ is a homeomorphism:
[…]
If the map $\mathscr{F}\rightarrow r_{*}r^{*}(\mathscr{F})$ is an isomorphism, then so is $g^{*}(\mathscr{F})\rightarrow g^{*}r_{*}r^{*}(\mathscr{F})$. Because $\overline{f}$ is flat, so is $g$, hence pullback along $g$ commutes with pushforward of quasi-coherent sheaves. Thus $g^{*}r_{*}r^{*}(\mathscr{F})=h_{*}t^{*}r^{*}(\mathscr{F})=h_{*}h^{*}(g^{*}(\mathscr{F}))$. So the map $g^{*}(\mathscr{F})\rightarrow h_{*}h^{*}(g^{*}(\mathscr{F}))$ is also an isomorphism. This proves the claim that whether a section of $j_{*}(\mathcal{O}_{Y\times_{X}Y})$ lies in $\mathscr{G}_{N}$ may be checked at codimension-one points.
Note that for any $u\in{\rm{H}}^{0}(\overline{Y}\times_{\overline{X}}\overline{Y},j_{*}\mathcal{O}_{Y\times_{X}Y})$, $u|Y\times_{X}Y$ lives in $g^{*}(\mathcal{O}(ND))|Y\times_{X}Y=\mathcal{O}$. Thus the only issue is whether $u$ lies in the image of $g^{*}(\mathcal{O}(ND))$ at the codimension-one points of the boundary of $\overline{Y}\times_{\overline{X}}\overline{Y}$. Because $\overline{Y}\times_{\overline{X}}\overline{Y}$ is a closed subscheme of $\overline{X}\otimes_{K}(L\otimes_{K}L)$, sections of the structure sheaf of $\overline{Y}\times_{\overline{X}}\overline{Y}$ all lie in the $(L\otimes_{K}L)$-span of sections of $\mathcal{O}_{\overline{X}}$, and for any section s of $\mathcal{O}_{\overline{X}}$, one has that $s\in\mathcal{O}_{\overline{X}}(nD)$ for some $n>0$ depending on $s$. It follows that any section of $\mathcal{O}_{\overline{Y}\times_{\overline{X}}\overline{Y}}$ lies in $\mathscr{G}_{n}$ for some $n>0$. We must show that there is a uniform bound on the $n$ required as $u$ varies over sections satisfying (4.2). Because $W$ is finite-dimensional, one has $W\subset\mathscr{G}_{m}$ for some $m>0$.
Let $z$ be one the finitely many boundary points of $\overline{Y}\times_{\overline{X}}\overline{Y}$. We will show that there is $N>0$ depending on $z$ such that every $u$ satisfying (4.2) lies in $\mathscr{G}_{N}$ near $z$. Since there are only finitely many $z$, this will suffice. Let $x:=g(z)$ be the image of $z$ in $\overline{X}$, and let $y\in\overline{Y}$ be the (unique) boundary point lying above $x$. Also let $B:=\mathcal{O}_{\overline{Y},y}$, $A:=\mathcal{O}_{\overline{X},x}$. Then $A$ is a Japanese – excellent even – DVR, and $B$ is finite flat over $A$. Furthermore, because $Y=(X_{L})_{\rm{red}}$, $B$ is spanned over $A$ by $L$. Thus $B^{p}\subset A$. We are therefore in the situation of Lemma 4.4. We claim that we may take $N=m+r-1$, where $r$ is as in that lemma.
To see this, let $\varpi$ be a uniformizer of $A$. Let $u$ satisfy (4.2), and let $n$ be the minimal nonnegative integer such that $u\in\mathscr{G}_{n}$ near $z$. We may assume that $n>0$ (as otherwise of course $u\in\mathscr{G}_{m+r-1}$ near $z$). We abuse notation and also denote by $u$ its image in $(B\otimes_{A}B)[1/\varpi]$. Because $u\in\mathscr{G}_{n}$, one has $\varpi^{n}u\in B\otimes_{A}B$. We claim that $\varpi^{n}u\notin\varpi(B\otimes_{A}B)$. To see this, note first that $B\otimes_{A}B$ – as a flat $A$-module – is $\varpi$-torsion free. So if $\varpi^{n}u\in\varpi(B\otimes_{A}B)$, then one would have $\varpi^{n-1}u\in B\otimes_{A}B$, hence $u\in\mathscr{G}_{n-1}$ near $z$, in violation of the minimality of $n$. Now suppose for the sake of contradiction that $n\geq m+r$, and multiply both sides of (4.2) by $\varpi^{2n}$. Because $W\subset\mathscr{G}_{m}$, we conclude that
$\pi_{12}^{*}(\varpi^{n}u)\pi_{23}^{*}(\varpi^{n}u)\in\varpi^{n-m}(B\otimes_{A}B\otimes_{A}B).$
Because $n-m\geq r$, this violates Proposition 4.2, applied with $I=I^{\prime}=(\varpi)$ and $I^{\prime\prime}=(\varpi^{r})$. We conclude, therefore, that $n\leq m+r-1$, as claimed. This completes the proof of the proposition. ∎
5 Almost-representability of the restricted Picard functor
In this section we prove the almost-representability of the restricted Picard functor for regular $K$-schemes of finite type. As a dévissage step, we show that, in proving Theorem 1.3, we are free to restrict to an open subscheme.
Lemma 5.1.
Let $K$ be a field, $X$ a finite type regular $K$-scheme and $U\subset X$ a dense open subscheme. Then $\operatorname{Pic}^{+}_{X/K}$ is almost-representable if and only if $\operatorname{Pic}^{+}_{U/K}$ is.
Proof.
Let $j\colon U\rightarrow X$ denote the inclusion, and consider the restriction map $j^{*}\colon\operatorname{Pic}^{+}_{X/K}\rightarrow\operatorname{Pic}^{+}_{U/K}$. For every smooth $K$-scheme $T$, $X\times T$ is regular, hence $\operatorname{Pic}(X\times T)\twoheadrightarrow\operatorname{Pic}(U\times T)$. It follows that $j^{*}$ is a surjection of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$. We claim that $\ker(j^{*})$ is étale with finitely generated group of $K_{s}$-points. By Galois descent, it suffices to show this after extending scalars to some finite Galois extension $L/K$, so we may assume that all of the codimension-$1$ components of $X\backslash U$ are geometrically irreducible. In that case, if the components are $D_{1},\dots,D_{n}$, then the natural map $\psi\colon\oplus_{D}\mathbf{Z}\rightarrow\ker(j^{*})$ is surjective. By Lemma 2.13, $\ker(\psi)$ is a constant group scheme, so $\ker(j^{*})$ is étale (with finitely generated group of $K_{s}$-points). The only if direction of the lemma now follows from Lemma 2.9.
Now we turn to the proof of the if direction (so $K$ need no longer be separably closed). Write $\operatorname{Pic}^{+}_{U/K}\simeq G/E$ with $G$ a smooth commutative $K$-group scheme with finitely generated component group, and $E\subset G^{t}$tale with $E(K_{s})$ free of finite rank, and let $H$ be defined by the following Cartesian square:
[…]
Then $H$ fits into an exact sequence
$0\longrightarrow\ker(j^{*})\longrightarrow H\longrightarrow G\longrightarrow 0$
Lemma 2.10 implies that $H$ is a (smooth) scheme, hence a smooth commutative $K$-group scheme. Its component group is finitely generated, because the same holds for $G$ and $\ker(j^{*})$. We have $\operatorname{Pic}^{+}_{X/K}\simeq H/E$, and Lemma 2.9 implies that $\operatorname{Pic}^{+}_{X/K}$ is almost-representable. ∎
Next we show that almost-representability descends through finite purely inseparable extensions.
Lemma 5.2.
Let $L/K$ be a finite purely inseparable extension field, let $X$ be a regular $K$-scheme of finite type, and suppose that $Y:=(X_{L})_{\rm{red}}$ is regular. If $\operatorname{Pic}^{+}_{Y/L}$ is almost-representable, then so is $\operatorname{Pic}^{+}_{X/K}$.
Proof.
[…]
We now prove the almost-representability of the restricted Picard functor.
Theorem 5.3.
For any regular $K$-scheme $X$ of finite type, $\operatorname{Pic}^{+}_{X/K}$ is almost-representable.
Proof.
If $\operatorname{Pic}^{+}_{X_{L}/L}$ is almost-representable for some finite Galois extension $L/K$, then so is $\operatorname{Pic}^{+}_{X_{L}/K}=\operatorname{R}_{L/K}(\operatorname{Pic}^{+}_{X_{L}/L})$, hence, by Proposition 2.15, so is $\operatorname{Pic}^{+}_{X/K}$. We may therefore assume that each component of $X$ is geometrically irreducible, and therefore, working component by component, that $X$ is geometrically irreducible. Then by [deJ2, Th. 5.9], there is a generically Galois alteration $Y\rightarrow X_{K_{\rm{perf}}}$ such that $\overline{Y}$ admits a smooth compactification. (That theorem applies when ${\rm{dim}}(X)>0$, but of course the 00-dimensional case is trivial.) The data descends from $K_{\rm{perf}}$ down to some finite purely inseparable extension $L/K$. Thus, there exist a finite purely inseparable extension $L/K$ and a proper, dominant, generically Galois morphism of $f\colon Y\rightarrow(X_{L})_{\rm{red}}$ from an integral $L$-scheme $Y$ such that $Y$ admits an $L$-smooth compactification $\overline{Y}$. Shrinking $X$ using lemma 5.1, we may assume that $f$ is finite Galois and that $(X_{L})_{\rm{red}}$ is regular.
The Picard functor $\operatorname{Pic}_{\overline{Y}/L}$ of the proper $L$-scheme $\overline{Y}$ is represented by a locally finite type $L$-scheme, which by the classical theorem of the base has finitely generated component group. Furthermore, this functor is the étale (rather than merely fppf) sheafification of the naive Picard functor $T\mapsto\operatorname{Pic}(\overline{Y}\times_{L}T)$ on the category of $L$-schemes, because $\overline{Y}$ is geometrically reduced. When we restrict to the category of smooth $L$-schemes, $\operatorname{Pic}_{\overline{Y}/L}$ agrees with the functor represented by its maximal smooth $L$-subgroup scheme. It follows that $\operatorname{Pic}^{+}_{\overline{Y}/L}$ is almost-representable (representable even). By Lemma 5.1, so too is $\operatorname{Pic}^{+}_{Y/L}$. Proposition 2.15 then implies that $\operatorname{Pic}^{+}_{(X_{L})_{\rm{red}}/L}$ is almost-representable, and Lemma 5.2 then implies the same for $\operatorname{Pic}^{+}_{X/K}$. ∎
Next we prove the uniqueness of the connected group $G^{0}$ arising in a presentation $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of $\operatorname{Pic}^{+}_{X/K}$ as an almost-representable functor.
Lemma 5.4.
Let $G$ be a normal connected commutative $K$-group scheme of finite type. Then $\operatorname{Ext}^{1}(G,\mathbf{Z}^{n})=0$.
Proof.
We may assume that $n=1$. Because $G$ is normal, ${\rm{H}}^{1}(G,\mathbf{Z})=0$. By [DG, Ch. II, §3, Prop. 2.3], therefore, the commutative extensions of $G$ by $\mathbf{Z}$ may be canonically identified with a subgroup of the second Hochschild cohomology group ${\rm{H}}^{2}_{0}(G,\mathbf{Z})$. Because $G$ is geometrically connected, $G\times G$ is connected, so any $2$-cocycle $G\times G\rightarrow\mathbf{Z}$ is constant, hence is a coboundary. It follows that $\operatorname{Ext}^{1}(G,\mathbf{Z})=0$. ∎
Lemma 5.5.
Let $G$ be a smooth connected $K$-group scheme, $\mathscr{F}$ a sheaf of abelian groups on the smooth-étale site of ${\rm{Spec}}(K)$, $E\subset\mathscr{F}$ a subsheaf represented by an étale $K$-group scheme such that $E(K_{s})$ is free of finite rank, and $\phi\colon G\rightarrow\mathscr{F}/E$ a homomorphism. Then $\phi$ lifts uniquely to a homomorphism $\psi\colon G\rightarrow\mathscr{F}$.
Proof.
Thanks to Galois descent and the uniqueness aspect of the lemma, we may assume that $E\simeq\mathbf{Z}^{n}$. The obstruction to lifting $\phi$ lies in $\operatorname{Ext}^{1}(G,E)$, which vanishes by Lemma 5.4. Given two lifts $\psi$, $\psi^{\prime}$, their difference yields a homomorphism $G\rightarrow E$, which must be trivial because $G$ is connected and $E$ is étale. ∎
We may now prove the uniqueness of $G^{0}$ in Theorem 1.3.
Proposition 5.6.
Let $G_{1},G_{2}$ be smooth $K$-group schemes, $E_{i}\subset G_{i}$ subsheaves on the smooth-étale site represented by étale $K$-group schemes such that $E_{i}(K_{s})$ is free of finite rank. Given an isomorphism $\phi\colon G_{1}/E_{1}\xrightarrow{\sim}G_{2}/E_{2}$, there is a unique map $\psi\colon G_{1}^{0}\rightarrow G_{2}^{0}$ such that the induced map $G_{1}^{0}/(E_{1}\cap G_{1}^{0})\rightarrow G_{2}^{0}/(E_{2}\cap G_{2}^{0})$ is the restriction of $\phi$. Further, $\psi$ is an isomorphism.
Proof.
Consider the map $G_{1}^{0}\rightarrow G_{1}/E_{1}\xrightarrow{\phi}G_{2}/E_{2}$. By Lemma 5.5, this lifts uniquely to a map $G_{1}^{0}\rightarrow G_{2}$, which must land inside $G_{2}^{0}$. This provides our unique $\psi$. To see that it is an isomorphism, apply the existence result to $\phi^{-1}$ to obtain a map $\gamma\colon G_{2}^{0}\rightarrow G_{1}^{0}$. Then $\gamma\circ\psi$ provides a lift on the connected component of the identity map of $G_{1}/E_{1}$, so by uniqueness it is the identity. A similar argument shows the same for $\psi\circ\gamma$. ∎
We now restate and prove the main Theorem 1.3.
Theorem 5.7.
((((Theorem 1.3)))) If $K$ is a field and $X$ is a regular $K$-scheme of finite type, then we have an isomorphism $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ of sheaves on the smooth-étale site of ${\rm{Spec}}(K)$, where $G$ is a smooth commutative $K$-group scheme with finitely generated component group and $E\subset G^{t}$ is an étale $K$-group scheme such $E(K_{s})$ is a free abelian group of finite rank. Furthermore, $G^{0}$ is unique up to unique isomorphism.
Proof.
Combine Theorem 5.3 and Proposition 5.6. ∎
6 Structure of the restricted Picard functor
In this section we investigate the structure of the restricted Picard functor, which – thanks to Theorem 5.3 – we now know to be almost-representable for regular, finite type $X/K$. In order to do this, we first require a means of determining when an element of $\operatorname{Pic}^{+}_{X/K}(S)$ arises from an element of $\operatorname{Pic}(X\times_{K}S)$, which is the purpose of the following proposition.
Proposition 6.1.
Let $X$ be a $K$-scheme, and $S$ a smooth $K$-scheme. Then there is an exact sequence, functorial in both $S$ and $X$:
$0\longrightarrow{\rm{H}}^{1}(S,f_{*}({\mathbf{G}}_{\rm{m}}))\longrightarrow\operatorname{Pic}(X\times_{K}S)\longrightarrow\operatorname{Pic}^{+}_{X/K}(S)\longrightarrow{\rm{H}}^{2}(S,f_{*}({\mathbf{G}}_{\rm{m}})),$
where $f\colon X\times S\rightarrow S$ is the projection, and the cohomology is étale. If $X(S)\neq\emptyset$, then we have an exact sequence
$\operatorname{Pic}(X\times_{K}S)\longrightarrow\operatorname{Pic}^{+}_{X/K}(S)\longrightarrow{\rm{H}}^{2}(S,f_{*}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}).$
Proof.
All terms in this proof are computed on the smooth-étale site of ${\rm{Spec}}(K)$. We first claim that $\operatorname{Pic}^{+}_{X/K}=\operatorname{R}^{1}f_{*}({\mathbf{G}}_{\rm{m}})$. Indeed, both sides are the sheafification of the presheaf $T\mapsto\operatorname{Pic}(X\times T)$. The first assertion of the proposition now follows from the Leray spectral sequence (which is functorial in both arguments)
$E_{2}^{i,j}={\rm{H}}^{i}(S,\operatorname{R}^{j}f_{*}({\mathbf{G}}_{\rm{m}}))\Longrightarrow{\rm{H}}^{i+j}(X\times S,{\mathbf{G}}_{\rm{m}}).$
For the second part, it is enough thanks to the spectral sequence to show that the map ${\rm{H}}^{2}(S,{\mathbf{G}}_{\rm{m}})\rightarrow{\rm{H}}^{2}(X\times S,{\mathbf{G}}_{\rm{m}})$ is injective, and this follows from the assumed existence of a section to $f$. ∎
Corollary 6.2.
If $X$ is a normal $K$-scheme of finite type such that each component admits a $K$-point, then there is a commutative étale $K$-group scheme $E$ such that $E(K_{s})$ is free of finite rank, and such that we have, for each smooth $K$-scheme $S$ an exact sequence, functorial in $X$ and $S$,
$\operatorname{Pic}(X\times_{K}S)\longrightarrow\operatorname{Pic}^{+}_{X/K}(S)\longrightarrow{\rm{H}}^{2}(S,E).$
Proof.
We may assume that $X$ is connected. Let $f\colon X\rightarrow{\rm{Spec}}(K)$ be the structure map. By Proposition 2.1, $f_{*}({\mathbf{G}}_{\rm{m}})$ – as a sheaf on the smooth-étale site of $K$ – lives in an exact sequence
$0\longrightarrow\operatorname{R}_{A/K}({\mathbf{G}}_{\rm{m}})\longrightarrow f_{*}({\mathbf{G}}_{\rm{m}})\longrightarrow E\longrightarrow 0,$
where $E$ is as in the corollary and $A\subset\Gamma(X,\mathcal{O}_{X})$ is the (finite) $K$-subalgebra of elements algebraic over $K$. By Proposition 6.1, therefore, we only need to show that $A=K$. Because $X$ is reduced and connected, $A$ is a field, and because $X(K)\neq\emptyset$, $A=K$. ∎
Lemma 6.3.
We have functorial isomorphisms, for every normal Noetherian scheme $S$, and every $i>0$, ${\rm{H}}^{i}(S,\mathbf{Q}/\mathbf{Z})\simeq{\rm{H}}^{i+1}(S,\mathbf{Z})$.
Proof.
This is a consequence of [Ros1, Lem. 5.2.9] and the exact sequence
$0\longrightarrow\mathbf{Z}\longrightarrow\mathbf{Q}\longrightarrow\mathbf{Q}/\mathbf{Z}\longrightarrow 0.\qed$
Our next goal is to show that $\operatorname{Pic}^{+}_{X/K}$ has no torus component and contains no copy of ${\mathbf{G}}_{\rm{a}}$ under mild hypotheses on $X$ (for instance, smoothness suffices). To this end, we prove the following proposition.
Proposition 6.4.
Let $X$ be a normal $K$-scheme of finite type such that each component of $X_{K_{s}}$ admits a $K_{s}$-point. Then there is no non-constant map $\mathbf{A}^{1}-\{0\}\rightarrow\operatorname{Pic}^{+}_{X/K}$.
Proof.
We may assume that $K$ is separably closed. Let $f\colon X\rightarrow{\rm{Spec}}(K)$ denote the structure map. Let $S:=\mathbf{A}^{1}-\{0\}$ for notational convenience. Suppose given an element $\alpha\in\operatorname{Pic}^{+}_{X/K}(S)$. By Corollary 6.2, the obstacle $\beta$ to $\alpha$ arising from a line bundle on $X\times S$ lives in ${\rm{H}}^{2}(S,\mathbf{Z}^{n})$ for some $n$. By Lemma 6.3, we have a functorial isomorphism ${\rm{H}}^{2}(S,\mathbf{Z})\simeq{\rm{H}}^{1}(S,\mathbf{Q}/\mathbf{Z})$. The latter agrees with ${\rm{Hom}}_{\rm{cts}}(\pi_{1}(S),\mathbf{Q}/\mathbf{Z})$. Now any finite connected étale cover of $S$ is given by a map $S\rightarrow S$. Passing to such a cover, therefore, we may assume that $\beta=0$, hence $\alpha$ arises from a line bundle on $X\times S$. Now $\operatorname{Pic}(X\times S)=\operatorname{Pic}(X)$ [EGAIV4, Cor. 21.4.11], so $\alpha$ actually arises from an element of $\operatorname{Pic}^{+}_{X/K}(K)$. That is, $\alpha$ is constant. ∎
Corollary 6.5.
Let $X$ be a normal $K$-scheme of finite type such that each component of $X_{K_{s}}$ admits a $K_{s}$-point, and let $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ be a presentation of $\operatorname{Pic}^{+}_{X/K}$ as an almost-representable functor. Then $G^{0}$ is an extension of a smooth unipotent group scheme by an abelian variety. If $K$ is perfect, then $G^{0}$ is an abelian variety.
Proof.
First suppose that $K$ is perfect. Then we may pass to the separable closure and assume that $K$ is algebraically closed. By Chevalley’s Theorem, $G^{0}$ is an extension of an abelian variety by a smooth connected commutative linear algebraic group $L$. Because $K$ is algebraically closed, $L={\mathbf{G}}_{\rm{m}}^{n}\times\mathbf{A}^{m}$ as a $K$-scheme for some $n,m\geq 0$. Proposition 6.4 therefore implies that $L=0$, so $G^{0}$ is an abelian variety.
Next suppose that $K$ has characteristic $p>0$. By [CGP, Th. A.3.9], one has an exact sequence of smooth connected $K$-group schemes
$0\longrightarrow A\longrightarrow G^{0}\longrightarrow L\longrightarrow 0$
with $A$ semi-abelian and $L$ affine. We claim that $A$ is an abelian variety and that $L$ is unipotent. We may assume that $K$ is separably closed. If $A$ were not an abelian variety, we would obtain a contradiction to Proposition 6.4. If $L$ were not unipotent, then it would contain a copy of ${\mathbf{G}}_{\rm{m}}$, hence we would obtain a subgroup $H$ of $G^{0}$ which is an extension of ${\mathbf{G}}_{\rm{m}}$ by an abelian variety. We claim that any such extension contains a copy of ${\mathbf{G}}_{\rm{m}}$, which would once again contradict Proposition 6.4.
To see the claim, note that $H$ defines an element of $\operatorname{Ext}^{1}({\mathbf{G}}_{\rm{m}},A)$. By Chevalley, over the perfect closure of $K$, $H$ must contain a copy of ${\mathbf{G}}_{\rm{m}}$ which necessarily maps isogenously onto the ${\mathbf{G}}_{\rm{m}}$ quotient. Because any endomorphism of ${\mathbf{G}}_{\rm{m}}$ over an extension field is already defined over $K$, we thus obtain a $K$-isogeny $\phi\colon{\mathbf{G}}_{\rm{m}}\rightarrow{\mathbf{G}}_{\rm{m}}$ such that $H$ splits upon pullback along $\phi$. Then the splitting composed with the map $\phi^{*}(H)\rightarrow H$ provides a nonzero homomorphism ${\mathbf{G}}_{\rm{m}}\rightarrow H$, whose image is then a copy of ${\mathbf{G}}_{\rm{m}}$ sitting inside $H$. ∎
Proposition 6.6.
If $X$ is a normal $K$-scheme of finite type such that $X(K)\neq\emptyset$ and $X_{K_{s}}$ has no non-constant units, then for every smooth $K$-scheme $S$, $\operatorname{Pic}^{+}_{X/K}(S)=\operatorname{Pic}(X\times S)/\operatorname{Pic}(S)$.
Proof.
The assumptions and Proposition 2.1 imply that $f_{*}({\mathbf{G}}_{\rm{m}})={\mathbf{G}}_{\rm{m}}$. The proposition therefore follows from Proposition 6.1. ∎
Proposition 6.7.
If $X$ is a smooth $K$-scheme of finite type such that each connected component of $X_{K_{s}}$ admits no non-constant units, then there is no non-constant map from a smooth unirational $K$-scheme to $\operatorname{Pic}^{+}_{X/K}$. In particular, there is no such map to $G$, where $G/E$ is a presentation of $\operatorname{Pic}^{+}_{X/K}$ as an almost-representable functor.
Proof.
We may assume that $K=K_{s}$ and that $X$ is nonempty and connected. Because $X$ is smooth, $X(K)\neq\emptyset$, so Proposition 6.6 implies that $\operatorname{Pic}(X\times S)\twoheadrightarrow\operatorname{Pic}^{+}_{X/K}(S)$ is surjective for every smooth $S/K$. We need to show that there are no nonconstant maps to $\operatorname{Pic}^{+}_{X/K}$ for $S$ unirational. We may assume that $S$ is open in $\mathbf{P}^{1}$. Then $\operatorname{Pic}^{+}_{X/K}(S)=\operatorname{Pic}(X\times S)=\operatorname{Pic}(X)$, the last equality by [EGAIV4, Cor. 21.4.11]. Thus, any map $S\rightarrow\operatorname{Pic}^{+}_{X/K}$ is constant – that is, is obtained via scalar extension from a map ${\rm{Spec}}(K)\rightarrow\operatorname{Pic}^{+}_{X/K}$. ∎
Our next goal is to obtain representability of $\operatorname{Pic}^{+}_{X/K}$ by a scheme when every map from $X$ into an abelian variety has finite image (Theorem 1.4).
Theorem 6.8.
((((Theorem 1.4)))) Let $K$ be a field and $X$ a regular $K$-scheme of finite type with dense smooth locus. Assume that any map from $X_{K_{s}}$ into an abelian variety $A/K_{s}$ factors through a finite subscheme of $A$. Then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative $K$-group scheme with finitely generated component group. Furthermore, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is wound unipotent.
Proof.
Let $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ be an almost-representable presentation of $\operatorname{Pic}^{+}_{X/K}$ (Theorem 5.3). By Corollary 6.5, $G^{0}$ is an extension of a smooth unipotent group scheme by an abelian variety $A$. The key point is to show that $A=0$. We may assume that $K$ is separably closed and that $X$ is nonempty and connected, hence irreducible. The map $A\rightarrow G/E\simeq\operatorname{Pic}^{+}_{X/K}$ yields an element $\alpha\in\operatorname{Pic}^{+}_{X/K}(A)$. By Corollary 6.2, the obstacle to $\alpha$ arising from an element of $\operatorname{Pic}(X\times A)$ lives in ${\rm{H}}^{2}(A,\mathbf{Z}^{n})$ for some $n\geq 0$. By Lemma 6.3, we have a functorial isomorphism ${\rm{H}}^{2}(A,\mathbf{Z})\simeq{\rm{H}}^{1}(A,\mathbf{Q}/\mathbf{Z})$. The latter agrees with ${\rm{Hom}}_{\rm{cts}}(\pi_{1}(A),\mathbf{Q}/\mathbf{Z})$. Any connected finite étale cover of $A$ is itself an abelian variety. We thus obtain an abelian variety $B$ – nonzero if $A$ is – and a finite homomorphism $B\rightarrow G^{0}$ such that the induced element of $\operatorname{Pic}^{+}_{X/K}(B)$ comes from a line bundle $\mathscr{L}$ on $X\times B$.
Now $\mathscr{L}$ defines a map $X\rightarrow\widehat{B}$ to the dual abelian variety. By hypothesis, this map has finite image. Because $X$ is integral and $X(K)\neq\emptyset$, it is a constant map to a point $b\in B(K)$. This means that $\mathscr{L}\in\operatorname{Pic}(X)\oplus\operatorname{Pic}(B)$, so the map $B\rightarrow\operatorname{Pic}^{+}_{X/K}$ is constant. This in turn means that the map $h\colon B\rightarrow G$ factors through $E\subset G$. Because $B$ is connected and $E$tale, it follows that $h=0$, hence $A=0$. By Corollary 6.5, it follows that $G^{0}$ is smooth and unipotent, and even trivial if $K$ is perfect. Thus $G^{t}$ is torsion. It follows that $E\subset G^{t}$ is trivial. Thus $\operatorname{Pic}^{+}_{X/K}\simeq G$ is a smooth $K$-scheme. Finally, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is wound unipotent by Proposition 6.4. ∎
Finally, we treat $\operatorname{Pic}^{+}_{X/K}$ when $X$ is a form of affine space (Theorem 1.6).
Theorem 6.9.
((((Theorem 1.6)))) If $K$ is a field and $X$ is a $K$-form of $\mathbf{A}^{n}_{K}$, then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative unipotent $K$-group scheme with strongly wound identity component. If $X(K)\neq\emptyset$, then, for every smooth $K$-scheme $S$, $\operatorname{Pic}^{+}_{X/K}(S)=\operatorname{Pic}(X\times S)/\operatorname{Pic}(S)$.
Proof.
The second statement of the theorem follows from Proposition 6.6, because $\mathbf{A}^{n}$ has no non-constant units over a field. Because $X$ is smooth, $X(K_{s})$ admits a $K_{s}$-point. Because $\mathbf{A}^{n}$ has no non-constant map to an abelian variety, Theorem 6.8 implies that $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth $K$-group scheme with finitely-generated component group. Proposition 6.7 implies that $(\operatorname{Pic}^{+}_{X/K})^{0}$ is strongly wound, so it only remains to show that the component group of $\operatorname{Pic}^{+}_{X/K}$ is $p$-power torsion, for which it suffices to show the same for $\operatorname{Pic}^{+}_{X/K}(K_{s})=\operatorname{Pic}(X_{K_{s}})$. Because $\operatorname{Pic}(X_{L})=\operatorname{Pic}(\mathbf{A}^{n}_{L})=0$ for some finite extension $L/K$, the claim follows from [Ros4, Lem. 3]. ∎
Example 6.10.
The conclusion that $\operatorname{Pic}^{+}_{X/K}(S)=\operatorname{Pic}(X\times S)/\operatorname{Pic}(S)$ in Theorem 6.9 really does require the assumption that $X(K)\neq\emptyset$ in general, even in the case $S={\rm{Spec}}(K)$, as the following example shows. (This also provides an example of a point of $\operatorname{Pic}^{+}_{X/K}(K)$ not arising from a line bundle without passing to an étale cover.) Denote by $f\colon X\rightarrow{\rm{Spec}}(K)$ the structure map. The spectral sequence
$E_{2}^{i,j}={\rm{H}}^{i}(K,\operatorname{R}^{j}f_{*}({\mathbf{G}}_{\rm{m}}))\Longrightarrow{\rm{H}}^{i+j}(X,{\mathbf{G}}_{\rm{m}})$
shows that the map $\operatorname{Pic}(X)\rightarrow\operatorname{Pic}^{+}_{X/K}(K)$ is surjective if and only if the pullback map ${\rm{H}}^{2}(K,{\mathbf{G}}_{\rm{m}})\rightarrow{\rm{H}}^{2}(X,{\mathbf{G}}_{\rm{m}})$ is injective. We will construct for any field $K$ of characteristic $p$ with $\operatorname{Br}(K)[p]\neq 0$ a $K$-form $X$ of $\mathbf{A}^{1}$ for which this injectivity fails.
By [Ros1, Prop. 2.6.4], for any regular ring $A$ of characteristic $p$, there is a functorial (in $A$) surjective group homomorphism
$\phi_{A}\colon\frac{\Omega^{1}_{A}}{dA+(C^{-1}-i)(\Omega^{1}_{A})}\twoheadrightarrow{\rm{H}}^{2}(A,{\mathbf{G}}_{\rm{m}})[p],$ (6.1)
where $d\colon A\rightarrow\Omega^{1}_{A}$ is the universal derivation, $C^{-1}\colon\Omega^{1}_{A}\rightarrow\Omega^{1}/dA$ is the inverse Cartier map defined by $C^{-1}(fdg)=f^{p}g^{p-1}dg$ (and extended by linearity) and $i\colon\Omega^{1}_{A}\rightarrow\Omega^{1}_{A}/dA$ is the projection. By assumption, ${\rm{H}}^{2}(K,{\mathbf{G}}_{\rm{m}})[p]\neq 0$, so it follows that there exist $\lambda,t\in K$ such that $\phi_{k}(\lambda dt)\neq 0\in\operatorname{Br}(K)$. But we will construct a $K$-form $X$ of $\mathbf{A}^{1}$ such that $\phi_{K}(\lambda dt)$ pulls back to $0\in{\rm{H}}^{2}(X,{\mathbf{G}}_{\rm{m}})$.
We define
$X:=\{Y^{p}+\lambda=t^{p-1}X^{p}-X\}\subset\mathbf{A}^{2}_{K}.$
First we check that $X$ is indeed a form of $\mathbf{A}^{1}$. Over $K(t^{1/p})$, one has the inverse isomorphisms
${X_{K(t^{1/p})}}$${\mathbf{A}^{1}_{K(t^{1/p})}}$${(X,Y)}$${Y-t^{(p-1)/p}X}$${(-S^{p}-\lambda,S-t^{(p-1)/p}S^{p}-t^{(p-1)/p}\lambda)}$${S}$
Now we check that $\phi_{K}(\lambda dt)\mapsto 0\in{\rm{H}}^{2}(X,{\mathbf{G}}_{\rm{m}})$. Indeed, if $B$ denotes the coordinate ring of $X$, then we have an equality in $\Omega^{1}_{B}/dB$,
$(C^{-1}-i)(Xdt)=(X^{p}t^{p-1}-X)dt=(Y^{p}+\lambda)dt=d(Y^{p}t)+\lambda dt.$
so $\lambda dt\in(C^{-1}-i)(\Omega^{1}_{B})$ as an element of $\Omega^{1}_{B}/dB$. By (6.1), therefore, we see that $f^{*}(\phi_{K}(\lambda dt))=\phi_{B}(\lambda dt)=0$.